### v1

# Model Lakes

Abstract.
Given a set of deep learning models, it can be hard to find models appropriate to a task, understand the models, and characterize how models are different one from another. Currently, practitioners rely on manually-written documentation to understand and choose models. However, not all models have complete and reliable documentation. As the number of machine learning models increases, this issue of finding, differentiating, and understanding models is becoming more crucial. Inspired from research on data lakes, we introduce and define the concept of model lakes. We discuss fundamental research challenges in the management of large models. And we discuss what principled data management techniques can be brought to bear on the study of large model management.
1. Introduction
With the rise of AI Chatbots (ChatGPT (OpenAI, [n.d.]a), BARD (Google, [n.d.]a), etc.) and other AI-integrated apps (AI, [n.d.]a; Microsoft, [n.d.]; AI, [n.d.]b; Labs, [n.d.]b; Google, [n.d.]b; Chintagunta et al., 2021; Trummer, 2022) many organizations commit significant resources to developing Machine Learning Models, most of which are fine-tuned versions of popular models such as BERT (Devlin et al., 2019), BART (Lewis et al., 2020), GPT-3 (Brown et al., 2020), LLAMA-2 (Touvron et al., 2023), and more. Recently, OpenAI created a mechanism (OpenAI, [n.d.]b) to allow users to create custom version of ChatGPT. To support this proliferation, sharing, and reuse (finetuning and editing) of large models, many models are hosted on platforms to support the collaborative use and sharing of models such as HuggingFace (Face, 2023) and Kaggle (Kaggle, 2023).
In the data management community, we have done a lot of introspection about how machine learning and Large Language models will revolutionize data management (Halevy et al., 2023; Fernandez et al., 2023). It is true that for many important (and stubbornly hard) semantic data management problems such as data discovery and data wrangling, advanced machine learning models are out performing other approaches (Fan et al., 2023b, a; Dong et al., 2023; Li et al., 2020; Narayan et al., 2022). Exciting research challenges remain in applying these models in data management and some point out the lack of exploration of models beyond LLMs (Borisov et al., 2022). However, we do not believe this makes machine learning an existential threat to data management or a challenge to the the raison d’être of data management (Fernandez et al., 2023). Rather, we argue that the large model revolution poses some new and fascinating scientific challenges that data management researchers are, perhaps, uniquely qualified to tackle. In this vision paper, we pose the question of how can principled data management revolutionize the use and management of models, including important problems like model selection (see Example 1) and model provenance. While model management (Bernstein et al., 2000) has long been a core topic in data management, we have been slow to move beyond the management of schemas and mappings (Doan et al., 2012) to the management of large machine learning models. Today, the way in which we search for models is through 20th century keyword search over (usually manually specified) names or metadata. The way in which we understand the relationship between two models is quite primitive (for example, is one manually declared to be a fine-tuned version of another). And the study of model provenance or lineage is just emerging (Mei et al., 2022; Mu et al., 2023).
We draw on lessons learned from collaborative data sharing platforms, such as Data Hub (Bhardwaj et al., 2015), that led to the development of data lakes (Nargesian et al., 2019) as powerful repositories for the management of heterogeneous collections of data. We introduce model lakes, as a parallel to data lakes, and discuss how important innovations in data lakes, including data discovery, annotation, and version management, can (and should) be applied within model lakes and studied with the same vigor.
Example 1 ().
Model Selection Problem: Consider a situation where a user wants to find a model that can summarize a legal document and simplify it in a non-technical manner. On HuggingFace (as of Jan 2024), the user finds that there are around 489,000 models uploaded and around 1500 of them have the ‘summarization’ task tag. While there are filters (trending, most likes, most downloads, model name search, and more), the user finds it hard to choose which model they should actually use. There are various concerns that the user goes through as she scrolls through different model cards (a common semi-structured form of model documentation). Is this model aware of legal jargon? Is it good at summarizing and simplifying legal documents? Is this the latest version of the model? Was this model trained on legal texts and if so which texts? What are other models that are similar to this model? Are they also trained on the same or different legal texts?
2. Models
Consider a model, , which is trained with dataset, . This model has an algorithm (function), , where . Based on the training objective provided by , can be discriminative or generative in nature. In the case of discriminative models, would learn to differentiate and classify different kinds of data instances. In the generative case, can be a sampling process that produces a distribution of new data instances. For instance, text-conditional diffusion models (Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022) take a text input, and produce an output that contain pixel values that form an image that matches the text description. For large language models (Touvron et al., 2023; Zheng et al., 2023; Brown et al., 2020), the output can be generated text (which may be in the format of tabular or semi-structured data), that the model creates in response to an input prompt . In general, training data can be of various types or combinations — it can be text documents, images, audio files, etc., that may be labeled or unlabeled. The full algorithm would include both ’s architecture and a pipeline for training and testing it. could be a parameterized model with any architecture, such as a Convolutional Neural Network (Cun et al., 1990) or a Transformer (Vaswani et al., 2017) or a simpler model. Note that unlike others (Fernandez et al., 2023), in our work, we consider not just LLMs, but also the worker-bee models that may be discriminative or generative, and that have been driving data science innovation since before the ChatGPT era. Moreover, recent findings from a survey on deep learning and tabular data (Borisov et al., 2022) underscore the necessity of investigating novel architectures for adoption, including diffusion or hybrid classical machine learning and deep learning models. They stress this is particularly important for handling heterogeneous collections of tabular data with better predictive performance and computational complexity. By developing and advancing a semantic model management system, the exploration of these models can be conducted in a more informed and efficient manner.
3. State-of-the-art in Model Management
Model Repositories and Model Registries. The proliferation of models has lead to the emergence of model repositories such as Civitai (Inc., 2023), HuggingFace (Face, 2023), or Kaggle (Kaggle, 2023) that openly share models. Companies, such as AWS also provide model repository products (such as the Amazon Elastic Container Registry) that can be used to share models within an organization. These platforms let users search for models by browsing through catalogs or by doing keyword search over manually created model names (and sometimes metadata). Beyond the hosting and sharing of models, a model registry provides clear naming conventions (for models and versions) and standardized, structured metadata (examples include Neptune (Labs, [n.d.]a) and AWS’ Sagemaker Model Registry (Verma et al., 2022)) and is more of a comprehensive system that enables documentation of the entire lifespan of machine learning models.
Documenting Models. Models that are deployed are usually accompanied with documentation known as model cards (Mitchell et al., 2019). This report generally contains the following categories of information.
Model Details: Information about the model such as model date, version, type, training algorithms used and other resources to learn more about the model.
Intended Use: Use cases that were imagined while developing the model and related primary users.
Factors: Summary of model performance across relevant features (like gender in people-centric models), instrumentation (like recording instrument for audio-based models), and environment (like image lighting for image-based models).
Metrics: Details of model performance measures, decision thresholds, and approaches to uncertainty and variability.
Evaluation Data: Information on dataset(s) used for quantitative analyses, the motivation behind choosing the dataset(s) and related preprocessing methods.
Training Data: Similar to information present for evaluation data section, but on data used to train the models.
Quantitative Analyses: Results of evaluating a model based on the metrics, which may be broken-down based on factors mentioned earlier. This includes unitary, i.e., model performance with respect to each factor, and intersectional results, i.e., model performance with respect to the intersection of factors.
Ethical Considerations: Information on ethical concerns such as model usage of sensitive data, intention of using the model for human life-related decisions, potential ethical risks and harms.
Caveats and Recommendations: Details of concerns or suggestions for using the model that are not covered elsewhere.
Similar to datasheets (Gebru et al., 2021), model cards are designed to guide developers in documenting models in a structured way. Model cards can be and should be augmented with information more similar to nutritional labels (Stoyanovich and Howe, 2019) that also include information about fairness and bias in the data (models). They can be further enriched with lineage and security related documentation such as adversarial attack and related defense measure policies, which are outlined in FactSheets for AI services (Arnold et al., 2019).
Model Search and Discovery. In Example 1, we showcase a potential scenario for the Model Selection Problem. The current solution pipeline involves a user searching for a relevant model by naming specific models or by typing related keyworks like legal to find models that either have that word in their name or in their model card. In other words, the search relies on the model’s name and documentation. Hence, any sorting of the answer by relevance is just the relevance (prevalence) of the keywords and is not a semantic notion. Of course, this search may fail if the documentation is incomplete or inaccurate. The state-of-the-art in verifying the documentation of a model is also in its infancy.
Interpretability. Interpretability has been defined as the extraction of relevant knowledge from a model concerning relationships either contained in data or learned by a model (Murdoch et al., 2019). Interpretability is critical for model analytics to enable semantic analysis and understanding of models and their training data. Several works have surveyed techniques and research questions in this area (Singh et al., 2024; Luo and Specia, 2024; Zhao et al., 2024; Räuker et al., 2022; Doshi-Velez and Kim, 2017; Gilpin et al., 2018), which can be broadly categorized into the following areas: Local model explanations explain the sensitivity of individual output predictions to local changes in inputs using gradients (Sundararajan et al., 2017), masks (Petsiuk et al., 2018; Fong et al., 2019), local models (Ribeiro et al., 2016), or Shapley values (Lundberg and Lee, 2017; Chen et al., 2023). Global explanations explain the mechanisms of the overall model at the level of attention patterns (Clark et al., 2019), representations (Belinkov, 2022; Zou et al., 2023), circuits (Elhage et al., 2021; Wang et al., 2022), or neurons (Bau et al., 2018, 2017; Geva et al., 2023). Models can be designed to be inherently interpretable (Rudin et al., 2022). Lastly, datasets can be explained using natural language explanations (Narayan et al., 2022), data visualizations (Dibia, 2023), or by training inherently interpretable models (Ustun and Rudin, 2016; Singh et al., 2023) of the data set.
Provenance. A new and important line of inquiry considers questions related to provenance (Mei et al., 2022; Mu et al., 2023). These issues pose similar questions to those studied in the database community (Buneman and Tan, 2018): from where did a generated fact derive or why was a predication made? A subset of these questions are also posed in model interpretability as part of local and global explanations. As in data provenance, the main issue is in whittling down all the provenance associated with some process (such as a model) into simple, but useful explanations (Buneman and Tan, 2018). More than two decades of research in the data management community have produced elegant and simple, yet powerful, models for data provenance when the process is a query (Buneman et al., 2008, 2016) or workflow (Moreau et al., 2013; Moreau and Groth, 2013). In large model provenance, the goal cannot be to understand and track all inputs (data, hyperparameters, code, training regimes, …) that were used to create a specific model output. Rather research is needed on finding meaningful sets of concepts that can be tracked efficiently and which provide important insights into model behavior.
There has been some initial research on solving the model provenance. Mu et al (Mu et al., 2023), for instance, uses a data-driven and model-driven representation learning method to encode ‘Model DNA’, which the authors then use as part of a framework for model provenance identification. Their goal is identify whether a model is a pre-trained model of another model. The former model is known as the source model while the latter model is known as the target model. This technique assumes a relaxed condition where the source and target model have the same architecture or structure and the target model’s training is based entirely on the source model. Furthermore, it assumes that we can access training data of the source model. However, there are harder conditions that can exist where the target model is trained based on a subset of the parameters of the source model or when the architecture of the target model is different from that of the source model. We define the granularity of such model conditions in Section 5 and explore various ways a model can be different from one another.
Model Versions. Models are valuable assets that are reused in numerous ways to create new versions. Based on the training type, the function or algorithm, can vary. For instance, if a model is pre-trained, would include a pipeline were it would encourage to learn general features from dataset . However, if is fine-tuned, would start off with weights from a previously trained model and learn features from to understand a specific task better. While pre-training and fine-tuning are some of the most common training objectives for creating new versions of a model, there are newer, more surgical-like training objectives such as model editing (Sinitsin et al., 2020; Meng et al., 2022) where focuses on updating certain facts (for instance, updating the name of the current President of a country) learned by a model in a more localized fashion. With various options for training methods, architecture, and dataset combinations, enormous numbers of models have been created with varying degrees of similarities and differences amongst each other.
Among the versioning methods, we examine fine-tuning and editing in more detail. Fine-tuning is the process of further training the parameters of an existing model to improve its performance, usually on specific task(s) or domain(s). For instance, T5 (Raffel et al., 2020) is a pre-trained model that has been trained on a collection of large collection of text (about 750 GB) to perform well on a diverse set of tasks. This has been further fine-tuned with the MIMIC-III (Johnson et al., 2016) and IV (Johnson et al., 2023) dataset to form Clinical-T5 (Lu et al., 2022) to perform better for medical domain-related tasks. Another model instance is FLAN-T5 (Chung et al., 2022) that has been fine-tuned on data with 1.8K tasks phrased as instruction so that it can handle such prompts better. Apart from datasets and training goals, there are also various optimizations schemes that could differentiate various fine-tuned models. For instance, Hu et al. (2022) experimented with a GPT-2 model fine-tuned with Adam (Kingma and Ba, 2015) and another GPT-2 model instance fine-tuned with their Low Rank Adaption (LoRA) (Hu et al., 2022). They both have the same task goals and are fine-tuned from the same source model, but have a difference in performance due to the fine-tuning optimization conducted. Consequently, with different datasets, training goals and algorithms, various versions of a pre-trained model can be created. Because they have the same source model, the fine-tuned models can also be considered to be related to each other.
Direct model Editing. Large language models (LLMs) contain a wealth of knowledge encoded in their parameters. As a result, they can respond to natural-language queries such as those that relate to factual statements. However, LLMs can sometimes hallucinate or point to outdated information. Hence, there have been several studies that look into localizing and understanding the internal mechanisms within LLMs. Zhu et al. (2020), for instance, proposed the task of modifying factual knowledge in transformer models such as BERT (Devlin et al., 2019) by performing a constrained fine-tuning on the original model with specific facts. De Cao et al. (2021), on the other hand, introduced Knowledge Editor, that uses hyper-networks (Ha et al., 2017) to tune the model’s parameters for editing specific knowledge. In terms of finding these facts, Dai et al. (2022) proposed the idea of knowledge neurons, whose activation highly correlated with their related knowledge expression. MEND (Mitchell et al., 2022) also edits local knowledge, but it does so by training a collection of Multi-Layer Perceptrons (MLP) to modify model gradients that would result in the production of local model edits. In contrast, ROME (Meng et al., 2022) uses rank one modification of the MLP weights of a single layer to directly edit appropriate weights that would change the fact associated to an information present the transformer model. MEMIT (Meng et al., 2023) builds upon ROME to be able to edit multiple knowledge expressions in the model. Direct model editing can also be applied to multimodal models: the TIME (Orgad et al., 2023) and UCE (Gandikota et al., 2024) methods edit text-to-image diffusion models. All these methods showcase various ways to edit a model, and hence, can be useful to understand how a current model is changed from its base model when a fact is changed.
4. Model Lakes: Vision
Given this state-of-the-art, we now outline our vision for model lakes: powerful repositories for the management of heterogeneous models. We enumerate challenges that require new scientific solutions and provide evidence that these challenges can be met.
Content-based Model Search. In analogy with dataset search, Google provides a comparable service to model registries, but for datasets (Benjelloun et al., 2020; Alrashed et al., 2021). They require dataset owners to follow their naming conventions and use metadata from Schema.org. Note this type of search only works well when there are clear incentives for artifact owners to provide complete and compliant metadata. Google provides this by returning datasets that comply with its best practices, though even Google struggles with impostors who game the system (Alrashed et al., 2021) by giving invalid semantic annotations. Current model repositories, such as Hugging Face, have found that they are subject to model designers who give sparse or ambiguous metadata within their model names or model cards (Liang et al., 2024). There is a danger that people could intentionally misinform with malicious intent (Security, [n.d.]b).
Within open data and enterprise data, researchers have learned that they cannot rely on metadata for datasets to be accurate, complete, or consistent within a data lake (Nargesian et al., 2019). As a result, there is a great deal of work on content-based search in data lakes including join search (Zhu et al., 2016, 2019; Khatiwada et al., 2022), union search (Nargesian et al., 2018; Khatiwada et al., 2023; Fan et al., 2023b), and related dataset search (Fernandez et al., 2018; Chapman et al., 2020). As mentioned in the introduction, machine learning models have revolutionized content-based dataset search. Important and impactful work has shown how we can use machine learning to create meaningful semantic representations of tuples, columns, and full tables to enhance dataset search and other semantic tasks like data integration and alignment.
But what about content-based model search? There is a bit of “physician, heal thyself” here. To the best of our knowledge, this is an area in its infancy. Lu et al. (2023) are the first to introduce the task of content-based model search, ”which aims to find the most relevant deep image generative models that satisfy a user’s input query”. Their solution builds a representation of the output distribution of a generative image model and uses this to find models that are most relevant to an input image or to an input model (or more precisely to its distribution). The trend of using meta-learning (and creating a hyper-network) for this task continues in other methods, such as AutoMRM, which is a model retrieval method based on multi-modal query and meta-learning (Li et al., 2023). A current limitation of these methods are that they are only able to perform retrieval for models that generate images. HuggingGPT (Shen et al., 2024) overcomes this limitation by using an LLM (in their case, ChatGPT) as a controller to decide which models to use based on a user’s prompt. This, however, is a metadata-based search and not a content based one, since the LLM would parse the user’s prompt into solvable tasks and then utilize another input of model descriptions (metadata) to decide which models would be relevant for each task. This approach allows queries and solutions in any modality or in any domain. However, it is limited by the capabilities of the LLM employed and the quality and accuracy of the model documentation. Furthermore, unlike content-based search, metadata-based ones would primarily fail when object-centric tags do not describe the model fully. As part of our vision, it is crucial that the content-based model search methods cover all types of models in the model lakes, including large language models. Additionally, ensuring their commercial usability, driven mainly by speed and accuracy, is essential. A hybrid approach combining metadata and content-based search could be effective, especially when exhaustive and verified metadata is accessible.
Related Model Search. Another important problem is related model search. A model is considered similar or related to another if they have some degree of commonality in one or more combinations of the following categories: datasets, training algorithms, behavior, or purpose. One approach to addressing related-model search was explored by Lu et al. (2023), which searches for image generative models by using the behavior of another image model as a query. We propose extending such model-query searches to all models in a model lake to help users identify related models when exploring the model in question. This can then be used as an informative analytics visualization as shown in Figure 1.
Documentation Verification and Auditing. It is crucial to be able to verify that manually created documentation is correct and complete. When documentation is missing or incomplete, can we learn it? This is similar to discovering and annotating metadata in data lakes (Sawadogo and Darmont, 2021; Zhao et al., 2021; Korini and Bizer, 2023; Fan et al., 2023a). If there is a documentation, the goal would be to verify them and add any information that was left out. Currently, Mithril AICert (Security, [n.d.]a), for instance, runs an initiative to provide certification to indicate that a model is trained with the specified algorithm and with the specified training dataset. However, there are a couple of caveats. First, this initiative is still under development and hence, cannot be used for production at the time of writing. Second, this mechanism is dependent on creators signing up to gain such certification. Furthermore, based on the limitations provided on AICert (Security, [n.d.]c) website, the training code and data itself cannot be audited for threats such as “backdoors injected into a model by the code or poisonous data.” Lastly, it also states that AICert also needs to be inspected since we are currently trusting that it is running at a safe server and does not have a malignant app from which users have to run their code on the training data. Hence, there are a couple of broad research questions to explore within the document-centric verification realm – What information do we need for a model to be reproducible? If the information is provided, how can we verify the information without being compromised in the process? If the information is not provided, how can we gather the necessary information from the models themselves? How can we automate this process? While developing models, there are systems like MISTIQUE (Vartak et al., 2018), that allows developers to perform model diagnosis, i.e., the process of examining machine learning model performance, and aiming to understand where the model excels and where it falls short. MISTIQUE can do so by capturing and storing the model intermediates, which are generally data artifacts that relate to the model including input data, prediction values, and data representations produced by model. These intermediate values can then be queried by the developers to diagnose the model to understand its performance. However, such frameworks are limited to white-box models, which are models where users can have full access including access to weights, activation, and gradients. However, we also need to consider documentation verification for other variations of model access such as black-box access, where users can only see the model’s output for their queries. Currently, there is no rigorous scheme present, especially for AI auditing (Casper et al., 2024).
Data Citation. Data Citation helps stakeholders identify the source, ownership, and authorship of the data used for a particular analysis. It is important to use proper data citation because the structure and contents of database can evolve (Buneman et al., 2016). Hence, this problem is also extended to machine learning research, since a big part of model training is its dataset. Based on a study conducted on 13K model cards (Liang et al., 2024), researchers found that if a training section of the model card is filled out, they generally include the name of the dataset, volume and characteristics of the dataset. This is not sufficient since the exact version of the dataset is usually not specified and the download links to the datasets may become stale. Furthermore, without any standardization and automated process on how to reference these datasets, many model cards might not contain such information or could feature varied statements regarding their referencing, as exemplified in the case shared by (Buneman et al., 2016).
In addition to citing datasets used during the model’s training and evaluation stages, we also need to consider how to manage data generated by these models. It is important to do so because users can use the generated information to train another model. One of the solutions for at least recognizing a generated data output is use watermarks (Kirchenbauer et al., 2023). Another scenario is when data is generated by a model and is being used by a consumer. It is important to know the model’s version and input prompts to the model to improve the chances of recreating this dataset. Hence, there is a need to advance data citation research for models to build an ecosystem where crucial information about related databases are recorded in both data and model lakes.
Provenance. A general research question for model provenance can be as follows: Given a potential source model, , and a single target model, , is the source model of ? If so, how is transformed into ? These transformations could be related to dataset being updated, training algorithms being changed, and/or changes in requirements of the model indicated by the developer as described by Mei et al. (2022). Inspired from data provenance, we can also explore where-, how- and why-provenance with respect to models. To get a comprehensive understanding of model provenance, it is important to know what data was used to train it. Using data citation helps us collect this information effectively. There are, however, alternative methods (Liu et al., 2023) that rely on intermediate model training checkpoints and not training data to at least verify model ownership. In terms of exploring why-provenance, we can try to extract datasets used or changed with the source and target models. For instance, there are a range of facts that can be edited using model editing techniques. A 1-fact edited model would mean that there was one fact used whose association is different from its source model. For instance, if ROME algorithm (Meng et al., 2022) was used to change the location of the Eiffel Tower from Paris to Rome, we intend to edit one sentence representation from “Eiffel Tower is located in Paris” to “Eiffel Tower is located in Rome”. Hence, the resulting model would be a 1-fact edited model of its source model. The first comparison that can be made within this sub-area, i.e., can we determine whether one model is 1-fact edited model of another. A general approach to understanding such models would be to reverse engineer the model’s prediction of outputs, which can lead to circuit discovery (Wang et al., 2022; Conmy et al., 2023; Elhage et al., 2021) or even understanding how to unlearn knowledge learned in models (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2023; Belrose et al., 2023; Kumari et al., 2023). We can extend the idea of such reverse engineering techniques to understand the process of 1-fact editing algorithms such as ROME on such models. Once such an ”edit or change” detector can be created, we could modify and extend its application. From model editing, we can then move on to model partial-tuning where a subset of the models’ parameters are used for training. We can look into the changed values for those parameters and understand what changed between its source and its current form. After building detectors for such cases, we can extend them further to understand the difference between one model that was fine-tuned from another model. Within fine-tuning, there are special cases of one-shot and few-shot training. Understanding these cases could help us build better detectors for the completely fine-tuned models where all model parameter values are usually changed. By creating such detectors, we can explain the transformation between a source and a target model, which is required for an exhaustive understanding of the target model’s provenance.
Model Version Search. While model provenance is a research challenge for identifying and explaining the transformation of one model, an extension to this challenge could be to identify versions given a set of models, and a set of datasets, . A model is considered a version of another when one can be transformed into another (such as through updates in dataset and/or training algorithm) and have the same true source model (i.e., the same first version). Apart from the approaches mentioned earlier, an alternative way to explore versioning is to understand how different entities have similar and/or different information present in various models. This could, for instance, be a post-hoc analysis through creation of knowledge graphs based on particular facts in data lakes that we know change from time to time, such as the President of the United States. By doing so, we could create a version history for data through models published at various time period, and by extension, identify the chronological order of the models based on their knowledge space.
Finally, by being able to pursue any direction indicated in the roadmap above, we can create a dataset or benchmark for such research so that the research community can build upon and create better methods for detecting such model differences. In addition to such benchmarks, it is also crucial to create metrics to understand how well these methods perform. By doing so, we can also increase outside-the-box access of AI models, where obtaining verified contextual information about the model can provide better transparency and evaluation for AI audits (Casper et al., 2024). Lastly, the solution to these methods would need visualization mechanisms to illustrate how these models and data are related. For instance, maybe we can have a hierarchical/network map of how different models are connected together? The next set of questions would be which parts of this mechanism can be automated and which ones would require a human-in-the-loop.
5. Granularity of Model Differences
There are various levels of granularity of model differences to consider. One aspect of granularity is explaining the transformation of a model. A low level granularity to identify the difference would be to identify the set of facts or data that were changed or were affected. A medium level of granularity could look like identifying the parameters that were changed and what dataset was it most likely trained on. A high level of granularity would be to just identify whether a model-in-question was fine-tuned from other model.
Another aspect of granularity is the level of information used to understand the model differences – did we use just the model weights? do we need their datasets? did we also use their documentation, i.e., their model cards? If we can find the difference between models by just using their model weights, then we can consider this difference to be a low-level information granularity. If we assume that we know at least one of the models’ datasets, then we can consider the difference to be a medium level. Lastly, if we depend on documentation to understand a pair of models’ differences, then we would consider it as a high-level of granularity. This can also scale the level of trust or confidence we can have for each model difference or similarity identification. For instance, identifying the difference between two models using just documentation has low level of confidence because we rely on trusting that the documentation is accurate. Furthermore, the accessibility of documentation can be limited because there is currently no structured way of documenting a model’s training process and inputs. However, if we rely on just the model’s weights, we can trust the identification more since it does not have an extra layer of information to process and understand the model.
Caveats and Considerations. There are a couple concerns with respect to extracting information about a model with low level granularity. If we were able to extract specific training data facts using just model weights, it can also be considered a privacy attack, which would mean that the model is unsafe in terms of privacy. Hence, in terms of providing granularity/verification scores as shown in Figure 1, it would have a high data verification score, but a low confidence in privacy. Another challenge with model lake tasks is differentiating between a predictive path and a deterministic path. For instance, in the case of identifying the source model of a certain model (model provenance task), we might find a couple of models that meet the possibility of it being a source model. In such cases, a lower than normal granularity score can be assigned unless a potential source model option is backed by another source of information, such as the model card.
6. MODEL LAKES: APPLICATION
Similar to a data lake, a model lake can be defined as a system or repository of many heterogeneous models and their related data stored in their natural format, for example, as files of model parameters. As the community progresses along the proposed directions, we can develop an intuitive navigation of a model lake, as depicted in Figure 1. This improvement can be helpful for addressing the Model Selection Problem shared in Example 1. In the left-side of Figure 1, we show that a user can search for models by typing out the task and type of data they are interested in working with. Based on this input, there can be an algorithm to semantically parse this text and find closely related task and data tags. Based on these tags, the search engine can share the most relevant models. Upon clicking a model, the model lake interface can lead to a model card page that may look similar to the right-side of the same figure. In this page, the users can view the model card description as well as what the model’s previous, next, source, and related models. In addition, they can also view what the training data was in order to understand if it was trained on desired text. If the training data is not specified, another way of understanding whether the model has a semantic understanding of the type of text the user cares about (for instance, legal jargon in the case of Example 1) is by viewing a semantic knowledge graph of the model and filtering for ‘legal’ within that graph. Within the model card, there can also be a privacy scale of low, medium, and high to gauge whether the model tends to share private information that was used in training data. Lastly, it could also have a set of verification scores (verifiable documentation task) to gauge how well this model is verified. The verification is categorized into data, code, and document verification. This scale would vary based on the level of granularity that they were verified with. 0 would indicate that they haven’t been verified and 1-5 would represent high level to low level verification where where high level implies less aspects of the model verified verses low level indicates that most, if not all, the verification was done using the model weight values.
7. Conclusion
The database community has responded to the “Big Model” revolution by proposing platforms like Agora (Traub et al., 2020), that manage data-related assets, including models, datasets, software, and compute resources in a coherent ecosystem. There is now a need for general methods for managing model lakes. It will be important that the methods generalize, irrespective of how many models we are trying to understand, how they are trained, or what base of information we wish to use to resolve our tasks. We call on the database community to contribute to the vision of model lakes, supporting users to better distinguish a set of models that they might want to choose from for their own tasks as well as alleviate future data and model management concerns that come up the rise of deep learning models. Our vision is for a fundamentally new platform that extends and integrates works on data/model provenance, data/model citation, and data/model version management.
Acknowledgements.
KP and RM was supported in part by NSF award numbers IIS-2107248, IIS-1956096, IIS-2325632, and KP and DB by a grant from Open Philanthropy.References
- (1)
- AI ([n.d.]a) ELSA AI. [n.d.]a. ELSA. https://elsaspeak.com/en/.
- AI ([n.d.]b) Fitness AI. [n.d.]b. Fitness AI. https://www.fitnessai.com/.
- Alrashed et al. (2021) Tarfah Alrashed, Dimitris Paparas, Omar Benjelloun, Ying Sheng, and Natasha F. Noy. 2021. Dataset or Not? A Study on the Veracity of Semantic Markup for Dataset Pages. In The Semantic Web - ISWC 2021 - 20th International Semantic Web Conference, ISWC 2021, Virtual Event, October 24-28, 2021, Proceedings (Lecture Notes in Computer Science), Andreas Hotho, Eva Blomqvist, Stefan Dietze, Achille Fokoue, Ying Ding, Payam M. Barnaghi, Armin Haller, Mauro Dragoni, and Harith Alani (Eds.), Vol. 12922. Springer, 338–356. https://doi.org/10.1007/978-3-030-88361-4_20
- Arnold et al. (2019) Matthew Arnold, Rachel KE Bellamy, Michael Hind, Stephanie Houde, Sameep Mehta, Aleksandra Mojsilović, Ravi Nair, K Natesan Ramamurthy, Alexandra Olteanu, David Piorkowski, et al. 2019. FactSheets: Increasing trust in AI services through supplier’s declarations of conformity. IBM Journal of Research and Development 63, 4/5 (2019), 6–1.
- Bau et al. (2018) Anthony Bau, Yonatan Belinkov, Hassan Sajjad, Nadir Durrani, Fahim Dalvi, and James Glass. 2018. Identifying and controlling important neurons in neural machine translation. arXiv preprint arXiv:1811.01157 (2018).
- Bau et al. (2017) David Bau, Bolei Zhou, Aditya Khosla, Aude Oliva, and Antonio Torralba. 2017. Network dissection: Quantifying interpretability of deep visual representations. In Proceedings of the IEEE conference on computer vision and pattern recognition. 6541–6549.
- Belinkov (2022) Yonatan Belinkov. 2022. Probing classifiers: Promises, shortcomings, and advances. Computational Linguistics 48, 1 (2022), 207–219.
- Belrose et al. (2023) Nora Belrose, David Schneider-Joseph, Shauli Ravfogel, Ryan Cotterell, Edward Raff, and Stella Biderman. 2023. LEACE: Perfect linear concept erasure in closed form. arXiv:2306.03819 [cs.LG]
- Benjelloun et al. (2020) Omar Benjelloun, Shiyu Chen, and Natasha F. Noy. 2020. Google Dataset Search by the Numbers. In The Semantic Web - ISWC 2020 - 19th International Semantic Web Conference, Athens, Greece, November 2-6, 2020, Proceedings, Part II (Lecture Notes in Computer Science), Jeff Z. Pan, Valentina A. M. Tamma, Claudia d’Amato, Krzysztof Janowicz, Bo Fu, Axel Polleres, Oshani Seneviratne, and Lalana Kagal (Eds.), Vol. 12507. Springer, 667–682. https://doi.org/10.1007/978-3-030-62466-8_41
- Bernstein et al. (2000) Philip A. Bernstein, Alon Y. Halevy, and Rachel Pottinger. 2000. A Vision of Management of Complex Models. SIGMOD Rec. 29, 4 (2000), 55–63. https://doi.org/10.1145/369275.369289
- Bhardwaj et al. (2015) Anant P. Bhardwaj, Amol Deshpande, Aaron J. Elmore, David R. Karger, Sam Madden, Aditya G. Parameswaran, Harihar Subramanyam, Eugene Wu, and Rebecca Zhang. 2015. Collaborative Data Analytics with DataHub. Proc. VLDB Endow. 8, 12 (2015), 1916–1919. https://doi.org/10.14778/2824032.2824100
- Borisov et al. (2022) Vadim Borisov, Tobias Leemann, Kathrin Seßler, Johannes Haug, Martin Pawelczyk, and Gjergji Kasneci. 2022. Deep Neural Networks and Tabular Data: A Survey. IEEE Transactions on Neural Networks and Learning Systems (2022), 1–21. https://doi.org/10.1109/TNNLS.2022.3229161
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems 33 (2020), 1877–1901.
- Buneman et al. (2008) Peter Buneman, James Cheney, and Stijn Vansummeren. 2008. On the expressiveness of implicit provenance in query and update languages. ACM Trans. Database Syst. 33, 4, Article 28 (dec 2008), 47 pages. https://doi.org/10.1145/1412331.1412340
- Buneman et al. (2016) Peter Buneman, Susan B. Davidson, and James Frew. 2016. Why data citation is a computational problem. Commun. ACM 59, 9 (2016), 50–57. https://doi.org/10.1145/2893181
- Buneman and Tan (2018) Peter Buneman and Wang-Chiew Tan. 2018. Data Provenance: What next? SIGMOD Rec. 47, 3 (2018), 5–16. https://doi.org/10.1145/3316416.3316418
- Casper et al. (2024) Stephen Casper, Carson Ezell, Charlotte Siegmann, Noam Kolt, Taylor Lynn Curtis, Benjamin Bucknall, Andreas Haupt, Kevin Wei, Jérémy Scheurer, Marius Hobbhahn, Lee Sharkey, Satyapriya Krishna, Marvin Von Hagen, Silas Alberti, Alan Chan, Qinyi Sun, Michael Gerovitch, David Bau, Max Tegmark, David Krueger, and Dylan Hadfield-Menell. 2024. Black-Box Access is Insufficient for Rigorous AI Audits. arXiv:2401.14446 [cs.CY]
- Chapman et al. (2020) Adriane Chapman, Elena Simperl, Laura Koesten, George Konstantinidis, Luis-Daniel Ibáñez, Emilia Kacprzak, and Paul Groth. 2020. Dataset search: a survey. VLDB J. 29, 1 (2020), 251–272. https://doi.org/10.1007/S00778-019-00564-X
- Chen et al. (2023) Hugh Chen, Ian C Covert, Scott M Lundberg, and Su-In Lee. 2023. Algorithms to estimate Shapley value feature attributions. Nature Machine Intelligence (2023), 1–12.
- Chintagunta et al. (2021) Bharath Chintagunta, Namit Katariya, Xavier Amatriain, and Anitha Kannan. 2021. Medically aware GPT-3 as a data generator for medical dialogue summarization. In Machine Learning for Healthcare Conference. PMLR, 354–372.
- Chung et al. (2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, Albert Webson, Shixiang Shane Gu, Zhuyun Dai, Mirac Suzgun, Xinyun Chen, Aakanksha Chowdhery, Alex Castro-Ros, Marie Pellat, Kevin Robinson, Dasha Valter, Sharan Narang, Gaurav Mishra, Adams Yu, Vincent Zhao, Yanping Huang, Andrew Dai, Hongkun Yu, Slav Petrov, Ed H. Chi, Jeff Dean, Jacob Devlin, Adam Roberts, Denny Zhou, Quoc V. Le, and Jason Wei. 2022. Scaling Instruction-Finetuned Language Models. arXiv:2210.11416 [cs.LG]
- Clark et al. (2019) Kevin Clark, Urvashi Khandelwal, Omer Levy, and Christopher D. Manning. 2019. What Does BERT Look at? An Analysis of BERT’s Attention. In Proceedings of the 2019 ACL Workshop BlackboxNLP: Analyzing and Interpreting Neural Networks for NLP, Tal Linzen, Grzegorz Chrupała, Yonatan Belinkov, and Dieuwke Hupkes (Eds.). Association for Computational Linguistics, Florence, Italy, 276–286. https://doi.org/10.18653/v1/W19-4828
- Conmy et al. (2023) Arthur Conmy, Augustine N. Mavor-Parker, Aengus Lynch, Stefan Heimersheim, and Adrià Garriga-Alonso. 2023. Towards Automated Circuit Discovery for Mechanistic Interpretability. arXiv:2304.14997 [cs.LG]
- Cun et al. (1990) Y. Le Cun, B. Boser, J. S. Denker, R. E. Howard, W. Habbard, L. D. Jackel, and D. Henderson. 1990. Handwritten digit recognition with a back-propagation network. Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 396–404.
- Dai et al. (2022) Damai Dai, Li Dong, Yaru Hao, Zhifang Sui, Baobao Chang, and Furu Wei. 2022. Knowledge Neurons in Pretrained Transformers. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). Association for Computational Linguistics, Dublin, Ireland, 8493–8502. https://doi.org/10.18653/v1/2022.acl-long.581
- De Cao et al. (2021) Nicola De Cao, Wilker Aziz, and Ivan Titov. 2021. Editing Factual Knowledge in Language Models. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, Online and Punta Cana, Dominican Republic, 6491–6506. https://doi.org/10.18653/v1/2021.emnlp-main.522
- Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers). Association for Computational Linguistics, Minneapolis, Minnesota, 4171–4186. https://doi.org/10.18653/v1/N19-1423
- Dibia (2023) Victor Dibia. 2023. LIDA: A Tool for Automatic Generation of Grammar-Agnostic Visualizations and Infographics using Large Language Models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics: System Demonstrations, ACL 2023, Toronto, Canada, July 10-12, 2023, Danushka Bollegala, Ruihong Huang, and Alan Ritter (Eds.). Association for Computational Linguistics, 113–126. https://doi.org/10.18653/V1/2023.ACL-DEMO.11
- Doan et al. (2012) AnHai Doan, Alon Y. Halevy, and Zachary G. Ives. 2012. Principles of Data Integration. Morgan Kaufmann. http://research.cs.wisc.edu/dibook/
- Dong et al. (2023) Yuyang Dong, Chuan Xiao, Takuma Nozawa, Masafumi Enomoto, and Masafumi Oyamada. 2023. DeepJoin: Joinable Table Discovery with Pre-trained Language Models. Proc. VLDB Endow. 16, 10 (2023), 2458 – 2470. https://doi.org/10.14778/3603581.3603587
- Doshi-Velez and Kim (2017) Finale Doshi-Velez and Been Kim. 2017. Towards a rigorous science of interpretable machine learning. arXiv preprint arXiv:1702.08608 (2017).
- Eldan and Russinovich (2023) Ronen Eldan and Mark Russinovich. 2023. Who’s Harry Potter? Approximate Unlearning in LLMs. arXiv:2310.02238 [cs.CL]
- Elhage et al. (2021) Nelson Elhage, Neel Nanda, Catherine Olsson, Tom Henighan, Nicholas Joseph, Ben Mann, Amanda Askell, Yuntao Bai, Anna Chen, Tom Conerly, Nova DasSarma, Dawn Drain, Deep Ganguli, Zac Hatfield-Dodds, Danny Hernandez, Andy Jones, Jackson Kernion, Liane Lovitt, Kamal Ndousse, Dario Amodei, Tom Brown, Jack Clark, Jared Kaplan, Sam McCandlish, and Chris Olah. 2021. A Mathematical Framework for Transformer Circuits. Transformer Circuits Thread (2021). https://transformer-circuits.pub/2021/framework/index.html.
- Face (2023) Hugging Face. 2023. https://huggingface.co
- Fan et al. (2023a) Grace Fan, Jin Wang, Yuliang Li, and Renée J. Miller. 2023a. Table Discovery in Data Lakes: State-of-the-art and Future Directions. In Companion of the 2023 International Conference on Management of Data (Seattle, WA, USA) (SIGMOD ’23). Association for Computing Machinery, New York, NY, USA, 69–75. https://doi.org/10.1145/3555041.3589409
- Fan et al. (2023b) Grace Fan, Jin Wang, Yuliang Li, Dan Zhang, and Renée J. Miller. 2023b. Semantics-aware Dataset Discovery from Data Lakes with Contextualized Column-based Representation Learning. PVDLB 16, 7 (2023), 1726–1739.
- Fernandez et al. (2018) Raul Castro Fernandez, Ziawasch Abedjan, Famien Koko, Gina Yuan, Samuel Madden, and Michael Stonebraker. 2018. Aurum: A Data Discovery System. In 34th IEEE International Conference on Data Engineering, ICDE 2018, Paris, France, April 16-19, 2018. IEEE Computer Society, 1001–1012. https://doi.org/10.1109/ICDE.2018.00094
- Fernandez et al. (2023) Raul Castro Fernandez, Aaron J. Elmore, Michael J. Franklin, Sanjay Krishnan, and Chenhao Tan. 2023. How Large Language Models Will Disrupt Data Management. Proc. VLDB Endow. 16, 11 (jul 2023), 3302–3309. https://doi.org/10.14778/3611479.3611527
- Fong et al. (2019) Ruth Fong, Mandela Patrick, and Andrea Vedaldi. 2019. Understanding deep networks via extremal perturbations and smooth masks. In Proceedings of the IEEE/CVF international conference on computer vision. 2950–2958.
- Gandikota et al. (2023) Rohit Gandikota, Joanna Materzyńska, Jaden Fiotto-Kaufman, and David Bau. 2023. Erasing Concepts from Diffusion Models. In Proceedings of the 2023 IEEE International Conference on Computer Vision.
- Gandikota et al. (2024) Rohit Gandikota, Hadas Orgad, Yonatan Belinkov, Joanna Materzyńska, and David Bau. 2024. Unified Concept Editing in Diffusion Models. IEEE/CVF Winter Conference on Applications of Computer Vision (2024).
- Gebru et al. (2021) Timnit Gebru, Jamie Morgenstern, Briana Vecchione, Jennifer Wortman Vaughan, Hanna Wallach, Hal Daumé Iii, and Kate Crawford. 2021. Datasheets for datasets. Commun. ACM 64, 12 (2021), 86–92.
- Geva et al. (2023) Mor Geva, Jasmijn Bastings, Katja Filippova, and Amir Globerson. 2023. Dissecting Recall of Factual Associations in Auto-Regressive Language Models. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 12216–12235. https://aclanthology.org/2023.emnlp-main.751
- Gilpin et al. (2018) Leilani H. Gilpin, David Bau, Ben Z. Yuan, Ayesha Bajwa, Michael A. Specter, and Lalana Kagal. 2018. Explaining Explanations: An Overview of Interpretability of Machine Learning. In 5th IEEE International Conference on Data Science and Advanced Analytics, DSAA 2018, Turin, Italy, October 1-3, 2018, Francesco Bonchi, Foster J. Provost, Tina Eliassi-Rad, Wei Wang, Ciro Cattuto, and Rayid Ghani (Eds.). IEEE, 80–89. https://doi.org/10.1109/DSAA.2018.00018
- Google ([n.d.]a) Google. [n.d.]a. BARD. https://bard.google.com/chat.
- Google ([n.d.]b) Google. [n.d.]b. Socratic. https://socratic.org/.
- Ha et al. (2017) David Ha, Andrew M. Dai, and Quoc V. Le. 2017. HyperNetworks. In International Conference on Learning Representations. https://openreview.net/forum?id=rkpACe1lx
- Halevy et al. (2023) Alon Y. Halevy, Yejin Choi, Avrilia Floratou, Michael J. Franklin, Natasha F. Noy, and Haixun Wang. 2023. Will LLMs reshape, supercharge, or kill data science? Proc. VLDB Endow. 16, 12 (2023), 4114–4115. https://doi.org/10.14778/3611540.3611634
- Hu et al. (2022) Edward J Hu, yelong shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2022. LoRA: Low-Rank Adaptation of Large Language Models. In International Conference on Learning Representations. https://openreview.net/forum?id=nZeVKeeFYf9
- Inc. (2023) Civit AI Inc. 2023. https://civitai.com/
- Jang et al. (2023) Joel Jang, Dongkeun Yoon, Sohee Yang, Sungmin Cha, Moontae Lee, Lajanugen Logeswaran, and Minjoon Seo. 2023. Knowledge Unlearning for Mitigating Privacy Risks in Language Models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 14389–14408. https://doi.org/10.18653/v1/2023.acl-long.805
- Johnson et al. (2023) Alistair EW Johnson, Lucas Bulgarelli, Lu Shen, Alvin Gayles, Ayad Shammout, Steven Horng, Tom J Pollard, Sicheng Hao, Benjamin Moody, Brian Gow, et al. 2023. MIMIC-IV, a freely accessible electronic health record dataset. Scientific data 10, 1 (2023), 1.
- Johnson et al. (2016) Alistair EW Johnson, Tom J Pollard, Lu Shen, Li-wei H Lehman, Mengling Feng, Mohammad Ghassemi, Benjamin Moody, Peter Szolovits, Leo Anthony Celi, and Roger G Mark. 2016. MIMIC-III, a freely accessible critical care database. Scientific data 3, 1 (2016), 1–9.
- Kaggle (2023) Kaggle. 2023. https://www.kaggle.com
- Khatiwada et al. (2023) Aamod Khatiwada, Grace Fan, Roee Shraga, Zixuan Chen, Wolfgang Gatterbauer, Renée J. Miller, and Mirek Riedewald. 2023. SANTOS: Relationship-based Semantic Table Union Search. In Accepted to appear in SIGMOD Conference. ACM. https://arxiv.org/pdf/2209.13589.pdf
- Khatiwada et al. (2022) Aamod Khatiwada, Roee Shraga, Wolfgang Gatterbauer, and Renée J Miller. 2022. Integrating Data Lake Tables. Proceedings of the VLDB Endowment 16, 4 (2022), 932–945.
- Kingma and Ba (2015) Diederik P. Kingma and Jimmy Ba. 2015. Adam: A Method for Stochastic Optimization. In 3rd International Conference on Learning Representations, ICLR 2015, San Diego, CA, USA, May 7-9, 2015, Conference Track Proceedings, Yoshua Bengio and Yann LeCun (Eds.). http://arxiv.org/abs/1412.6980
- Kirchenbauer et al. (2023) John Kirchenbauer, Jonas Geiping, Yuxin Wen, Jonathan Katz, Ian Miers, and Tom Goldstein. 2023. A Watermark for Large Language Models. In Proceedings of the 40th International Conference on Machine Learning (Proceedings of Machine Learning Research), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.), Vol. 202. PMLR, 17061–17084. https://proceedings.mlr.press/v202/kirchenbauer23a.html
- Korini and Bizer (2023) Keti Korini and Christian Bizer. 2023. Column Type Annotation using ChatGPT. In Joint Proceedings of Workshops at the 49th International Conference on Very Large Data Bases (VLDB 2023), Vancouver, Canada, August 28 - September 1, 2023 (CEUR Workshop Proceedings), Rajesh Bordawekar, Cinzia Cappiello, Vasilis Efthymiou, Lisa Ehrlinger, Vijay Gadepally, Sainyam Galhotra, Sandra Geisler, Sven Groppe, Le Gruenwald, Alon Y. Halevy, Hazar Harmouch, Oktie Hassanzadeh, Ihab F. Ilyas, Ernesto Jiménez-Ruiz, Sanjay Krishnan, Tirthankar Lahiri, Guoliang Li, Jiaheng Lu, Wolfgang Mauerer, Umar Farooq Minhas, Felix Naumann, M. Tamer Özsu, El Kindi Rezig, Kavitha Srinivas, Michael Stonebraker, Satyanarayana R. Valluri, Maria-Esther Vidal, Haixun Wang, Jiannan Wang, Yingjun Wu, Xun Xue, Mohamed Zaït, and Kai Zeng (Eds.), Vol. 3462. CEUR-WS.org. https://ceur-ws.org/Vol-3462/TADA1.pdf
- Kumari et al. (2023) Nupur Kumari, Bingliang Zhang, Sheng-Yu Wang, Eli Shechtman, Richard Zhang, and Jun-Yan Zhu. 2023. Ablating Concepts in Text-to-Image Diffusion Models. In International Conference on Computer Vision (ICCV).
- Labs ([n.d.]a) Neptune Labs. [n.d.]a. neptune.ai — The MLOps stack component for experiment tracking. https://neptune.ai/.
- Labs ([n.d.]b) Prisma Labs. [n.d.]b. Lensa. https://prisma-ai.com/lensa.
- Lewis et al. (2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020. BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. Association for Computational Linguistics, Online, 7871–7880. https://doi.org/10.18653/v1/2020.acl-main.703
- Li et al. (2020) Yuliang Li, Jinfeng Li, Yoshihiko Suhara, AnHai Doan, and Wang-Chiew Tan. 2020. Deep Entity Matching with Pre-Trained Language Models. 14, 1 (sep 2020), 50–60. https://doi.org/10.14778/3421424.3421431
- Li et al. (2023) Zhaotian Li, Binhang Qi, Hailong Sun, and Xiang Gao. 2023. AutoMRM: A Model Retrieval Method Based on Multimodal Query and Meta-learning (CIKM ’23). Association for Computing Machinery, New York, NY, USA, 1228–1237. https://doi.org/10.1145/3583780.3614787
- Liang et al. (2024) Weixin Liang, Nazneen Rajani, Xinyu Yang, Ezinwanne Ozoani, Eric Wu, Yiqun Chen, Daniel Scott Smith, and James Zou. 2024. What’s documented in AI? Systematic Analysis of 32K AI Model Cards. arXiv:2402.05160 [cs.SE]
- Liu et al. (2023) Yunpeng Liu, Kexin Li, Zhuotao Liu, Bihan Wen, Ke Xu, Weiqiang Wang, Wenbiao Zhao, and Qi Li. 2023. Provenance of Training without Training Data: Towards Privacy-Preserving DNN Model Ownership Verification. In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23). Association for Computing Machinery, New York, NY, USA, 1980–1990. https://doi.org/10.1145/3543507.3583198
- Lu et al. (2023) Daohan Lu, Sheng-Yu Wang, Nupur Kumari, Rohan Agarwal, Mia Tang, David Bau, and Jun-Yan Zhu. 2023. Content-based Search for Deep Generative Models. In SIGGRAPH Asia 2023 Conference Papers (SA ’23). ACM. https://doi.org/10.1145/3610548.3618189
- Lu et al. (2022) Qiuhao Lu, Dejing Dou, and Thien Nguyen. 2022. ClinicalT5: A Generative Language Model for Clinical Text. In Findings of the Association for Computational Linguistics: EMNLP 2022, Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang (Eds.). Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 5436–5443. https://doi.org/10.18653/v1/2022.findings-emnlp.398
- Lundberg and Lee (2017) Scott M. Lundberg and Su-In Lee. 2017. A unified approach to interpreting model predictions. CoRR abs/1705.07874 (2017). arXiv:1705.07874 http://arxiv.org/abs/1705.07874
- Luo and Specia (2024) Haoyan Luo and Lucia Specia. 2024. From Understanding to Utilization: A Survey on Explainability for Large Language Models. CoRR abs/2401.12874 (2024). https://doi.org/10.48550/ARXIV.2401.12874 arXiv:2401.12874
- Mei et al. (2022) Songzhu Mei, Cong Liu, Qinglin Wang, and Huayou Su. 2022. Model Provenance Management in MLOps Pipeline. In Proceedings of the 2022 8th International Conference on Computing and Data Engineering (Bangkok, Thailand) (ICCDE ’22). Association for Computing Machinery, New York, NY, USA, 45–50. https://doi.org/10.1145/3512850.3512861
- Meng et al. (2022) Kevin Meng, David Bau, Alex Andonian, and Yonatan Belinkov. 2022. Locating and Editing Factual Associations in GPT. Advances in Neural Information Processing Systems 36 (2022).
- Meng et al. (2023) Kevin Meng, Arnab Sen Sharma, Alex Andonian, Yonatan Belinkov, and David Bau. 2023. Mass Editing Memory in a Transformer. The Eleventh International Conference on Learning Representations (ICLR) (2023).
- Microsoft ([n.d.]) Microsoft. [n.d.]. Seeing AI. https://www.microsoft.com/en-us/ai/seeing-ai.
- Mitchell et al. (2022) Eric Mitchell, Charles Lin, Antoine Bosselut, Chelsea Finn, and Christopher D Manning. 2022. Fast Model Editing at Scale. In International Conference on Learning Representations. https://openreview.net/pdf?id=0DcZxeWfOPt
- Mitchell et al. (2019) Margaret Mitchell, Simone Wu, Andrew Zaldivar, Parker Barnes, Lucy Vasserman, Ben Hutchinson, Elena Spitzer, Inioluwa Deborah Raji, and Timnit Gebru. 2019. Model Cards for Model Reporting. In Proceedings of the Conference on Fairness, Accountability, and Transparency (Atlanta, GA, USA) (FAT* ’19). Association for Computing Machinery, New York, NY, USA, 220–229. https://doi.org/10.1145/3287560.3287596
- Moreau and Groth (2013) Luc Moreau and Paul Groth. 2013. Provenance: An Introduction to PROV. Morgan & Claypool Publishers. https://doi.org/10.2200/S00528ED1V01Y201308WBE007
- Moreau et al. (2013) Luc Moreau, Paolo Missier, Khalid Belhajjame, Reza B’Far, James Cheney, Sam Coppens, Stephen Cresswell, Yolanda Gil, Paul Groth, Graham Klyne, et al. 2013. Prov-dm: The prov data model.
- Mu et al. (2023) Xin Mu, Yu Wang, Yehong Zhang, Jiaqi Zhang, Hui Wang, Yang Xiang, and Yue Yu. 2023. Model Provenance via Model DNA. arXiv:2308.02121 [cs.LG]
- Murdoch et al. (2019) W. James Murdoch, Chandan Singh, Karl Kumbier, Reza Abbasi-Asl, and Bin Yu. 2019. Definitions, methods, and applications in interpretable machine learning. Proceedings of the National Academy of Sciences 116, 44 (2019), 22071–22080. https://doi.org/10.1073/pnas.1900654116 arXiv:https://www.pnas.org/doi/pdf/10.1073/pnas.1900654116
- Narayan et al. (2022) Avanika Narayan, Ines Chami, Laurel Orr, and Christopher Ré. 2022. Can Foundation Models Wrangle Your Data? Proc. VLDB Endow. 16, 4 (dec 2022), 738–746. https://doi.org/10.14778/3574245.3574258
- Nargesian et al. (2019) Fatemeh Nargesian, Erkang Zhu, Renée J. Miller, Ken Q. Pu, and Patricia C. Arocena. 2019. Data Lake Management: Challenges and Opportunities. PVLDB 12, 12 (2019), 1986–1989.
- Nargesian et al. (2018) Fatemeh Nargesian, Erkang Zhu, Ken Q Pu, and Renée J Miller. 2018. Table union search on open data. PVLDB 11, 7 (2018), 813–825.
- OpenAI ([n.d.]a) OpenAI. [n.d.]a. Chat GPT. https://chat.openai.com/chat.
- OpenAI ([n.d.]b) OpenAI. [n.d.]b. Introducing GPTs. https://openai.com/blog/introducing-gpts.
- Orgad et al. (2023) Hadas Orgad, Bahjat Kawar, and Yonatan Belinkov. 2023. Editing Implicit Assumptions in Text-to-Image Diffusion Models. In IEEE/CVF International Conference on Computer Vision, ICCV 2023, Paris, France, October 1-6, 2023. IEEE, 7030–7038. https://doi.org/10.1109/ICCV51070.2023.00649
- Petsiuk et al. (2018) Vitali Petsiuk, Abir Das, and Kate Saenko. 2018. Rise: Randomized input sampling for explanation of black-box models. arXiv preprint arXiv:1806.07421 (2018).
- Raffel et al. (2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J Liu. 2020. Exploring the limits of transfer learning with a unified text-to-text transformer. The Journal of Machine Learning Research 21, 1 (2020), 5485–5551.
- Ramesh et al. (2022) Aditya Ramesh, Prafulla Dhariwal, Alex Nichol, Casey Chu, and Mark Chen. 2022. Hierarchical Text-Conditional Image Generation with CLIP Latents. arXiv:2204.06125 [cs.CV]
- Räuker et al. (2022) Tilman Räuker, Anson Ho, Stephen Casper, and Dylan Hadfield-Menell. 2022. Toward Transparent AI: A Survey on Interpreting the Inner Structures of Deep Neural Networks. CoRR abs/2207.13243 (2022). https://doi.org/10.48550/ARXIV.2207.13243 arXiv:2207.13243
- Ribeiro et al. (2016) Marco Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. 2016. ” Why should i trust you?” Explaining the predictions of any classifier. In Proceedings of the 22nd ACM SIGKDD international conference on knowledge discovery and data mining. 1135–1144.
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-resolution image synthesis with latent diffusion models. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 10684–10695.
- Rudin et al. (2022) Cynthia Rudin, Chaofan Chen, Zhi Chen, Haiyang Huang, Lesia Semenova, and Chudi Zhong. 2022. Interpretable machine learning: Fundamental principles and 10 grand challenges. Statistic Surveys 16 (2022), 1–85.
- Saharia et al. (2022) Chitwan Saharia, William Chan, Saurabh Saxena, Lala Li, Jay Whang, Emily Denton, Seyed Kamyar Seyed Ghasemipour, Burcu Karagol Ayan, S. Sara Mahdavi, Rapha Gontijo Lopes, Tim Salimans, Jonathan Ho, David J Fleet, and Mohammad Norouzi. 2022. Photorealistic Text-to-Image Diffusion Models with Deep Language Understanding. arXiv:2205.11487 [cs.CV]
- Sawadogo and Darmont (2021) Pegdwendé Sawadogo and Jérôme Darmont. 2021. On data lake architectures and metadata management. Journal of Intelligent Information Systems 56 (2021), 97–120.
- Security ([n.d.]a) Mithril Security. [n.d.]a. AICert – Open-source tool to trace AI models’ provenance. https://www.mithrilsecurity.io/aicert.
- Security ([n.d.]b) Mithril Security. [n.d.]b. PoisonGPT: How We Hid a Lobotomized LLM on Hugging Face to Spread Fake News. https://blog.mithrilsecurity.io/poisongpt-how-we-hid-a-lobotomized-llm-on-hugging-face-to-spread-fake-news/.
- Security ([n.d.]c) Mithril Security. [n.d.]c. Welcome to AICert! https://aicert.mithrilsecurity.io/en/latest/.
- Shen et al. (2024) Yongliang Shen, Kaitao Song, Xu Tan, Dongsheng Li, Weiming Lu, and Yueting Zhuang. 2024. Hugginggpt: Solving ai tasks with chatgpt and its friends in hugging face. Advances in Neural Information Processing Systems 36 (2024).
- Singh et al. (2023) Chandan Singh, Armin Askari, Rich Caruana, and Jianfeng Gao. 2023. Augmenting interpretable models with large language models during training. Nature Communications 14, 1 (2023), 7913.
- Singh et al. (2024) Chandan Singh, Jeevana Priya Inala, Michel Galley, Rich Caruana, and Jianfeng Gao. 2024. Rethinking Interpretability in the Era of Large Language Models. arXiv:2402.01761 [cs.CL]
- Sinitsin et al. (2020) Anton Sinitsin, Vsevolod Plokhotnyuk, Dmitry Pyrkin, Sergei Popov, and Artem Babenko. 2020. Editable Neural Networks. In International Conference on Learning Representations. https://openreview.net/forum?id=HJedXaEtvS
- Stoyanovich and Howe (2019) Julia Stoyanovich and Bill Howe. 2019. Nutritional Labels for Data and Models. IEEE Data Eng. Bull. 42, 3 (2019), 13–23. http://sites.computer.org/debull/A19sept/p13.pdf
- Sundararajan et al. (2017) Mukund Sundararajan, Ankur Taly, and Qiqi Yan. 2017. Axiomatic attribution for deep networks. In International conference on machine learning. PMLR, 3319–3328.
- Touvron et al. (2023) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023. Llama 2: Open Foundation and Fine-Tuned Chat Models. arXiv:2307.09288 [cs.CL]
- Traub et al. (2020) Jonas Traub, Zoi Kaoudi, Jorge-Arnulfo Quiané-Ruiz, and Volker Markl. 2020. Agora: Bringing Together Datasets, Algorithms, Models and More in a Unified Ecosystem [Vision]. SIGMOD Rec. 49, 4 (2020), 6–11. https://doi.org/10.1145/3456859.3456861
- Trummer (2022) Immanuel Trummer. 2022. CodexDB: Synthesizing code for query processing from natural language instructions using GPT-3 Codex. Proceedings of the VLDB Endowment 15, 11 (2022), 2921–2928.
- Ustun and Rudin (2016) Berk Ustun and Cynthia Rudin. 2016. Supersparse linear integer models for optimized medical scoring systems. Machine Learning 102 (2016), 349–391.
- Vartak et al. (2018) Manasi Vartak, Joana M. F. da Trindade, Samuel Madden, and Matei Zaharia. 2018. MISTIQUE: A System to Store and Query Model Intermediates for Model Diagnosis. In Proceedings of the 2018 International Conference on Management of Data (Houston, TX, USA) (SIGMOD ’18). Association for Computing Machinery, New York, NY, USA, 1285–1300. https://doi.org/10.1145/3183713.3196934
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N. Gomez, Łukasz Kaiser, and Illia Polosukhin. 2017. Attention is all you need. In Proceedings of the 31st International Conference on Neural Information Processing Systems (Long Beach, California, USA) (NIPS’17). Curran Associates Inc., Red Hook, NY, USA, 6000–6010.
- Verma et al. (2022) Sandeep Verma, Farooq Sabir, Mani Khanuja, Rupinder Grewal, Saumitra Vikram, and Sreedevi Srinivasan. 2022. Build a cross-account MLOps workflow using the Amazon SageMaker model registry. https://aws.amazon.com/blogs/machine-learning/build-a-cross-account-mlops-workflow-using-the-amazon-sagemaker-model-registry/.
- Wang et al. (2022) Kevin Ro Wang, Alexandre Variengien, Arthur Conmy, Buck Shlegeris, and Jacob Steinhardt. 2022. Interpretability in the Wild: a Circuit for Indirect Object Identification in GPT-2 Small. In The Eleventh International Conference on Learning Representations.
- Zhao et al. (2024) Haiyan Zhao, Fan Yang, Himabindu Lakkaraju, and Mengnan Du. 2024. Opening the Black Box of Large Language Models: Two Views on Holistic Interpretability. CoRR abs/2402.10688 (2024). https://doi.org/10.48550/ARXIV.2402.10688 arXiv:2402.10688
- Zhao et al. (2021) Yan Zhao, Franck Ravat, Julien Aligon, Chantal Soule-dupuy, Gabriel Ferrettini, and Imen Megdiche. 2021. Analysis-oriented Metadata for Data Lakes. In Proceedings of the 25th International Database Engineering & Applications Symposium (Montreal, QC, Canada) (IDEAS ’21). Association for Computing Machinery, New York, NY, USA, 194–203. https://doi.org/10.1145/3472163.3472273
- Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric. P Xing, Hao Zhang, Joseph E. Gonzalez, and Ion Stoica. 2023. Judging LLM-as-a-judge with MT-Bench and Chatbot Arena. arXiv:2306.05685 [cs.CL]
- Zhu et al. (2020) Chen Zhu, Ankit Singh Rawat, Manzil Zaheer, Srinadh Bhojanapalli, Daliang Li, Felix Yu, and Sanjiv Kumar. 2020. Modifying Memories in Transformer Models. arXiv:2012.00363 [cs.CL]
- Zhu et al. (2019) Erkang Zhu, Dong Deng, Fatemeh Nargesian, and Renée J Miller. 2019. Josie: Overlap set similarity search for finding joinable tables in data lakes. In SIGMOD. 847–864.
- Zhu et al. (2016) Erkang Zhu, Fatemeh Nargesian, Ken Q. Pu, and Renée J. Miller. 2016. LSH Ensemble: Internet-Scale Domain Search. Proc. VLDB Endow. 9, 12 (2016), 1185–1196. https://doi.org/10.14778/2994509.2994534
- Zou et al. (2023) Andy Zou, Long Phan, Sarah Chen, James Campbell, Phillip Guo, Richard Ren, Alexander Pan, Xuwang Yin, Mantas Mazeika, Ann-Kathrin Dombrowski, et al. 2023. Representation engineering: A top-down approach to ai transparency. arXiv preprint arXiv:2310.01405 (2023).

### v2

@@
 However, not all models have complete and reliable documentation.
-As the number of machine learning models increases, this issue of finding, differentiating, and understanding models is becoming more crucial.
-Inspired from research on data lakes, we introduce and define the concept of model lakes.
-We discuss fundamental research challenges in the management of large models.
-And we discuss what principled data management techniques can be brought to bear on the study of large model management.
+As the number of models increases, the challenges of finding, differentiating, and understanding models become increasingly crucial.
+Inspired from research on data lakes, we introduce the concept of model lakes.
+We formalize key model lake tasks, including model attribution, versioning, search, and benchmarking, and discuss fundamental research challenges in the management of large models.
+We also explore what data management techniques can be brought to bear on the study of large model management.
 1. Introduction
-With the rise of AI Chatbots (ChatGPT (OpenAI, [n.d.]a), BARD (Google, [n.d.]a), etc.) and other AI-integrated apps (AI, [n.d.]a; Microsoft, [n.d.]; AI, [n.d.]b; Labs, [n.d.]b; Google, [n.d.]b; Chintagunta et al., 2021; Trummer, 2022) many organizations commit significant resources to developing Machine Learning Models, most of which are fine-tuned versions of popular models such as BERT (Devlin et al., 2019), BART (Lewis et al., 2020), GPT-3 (Brown et al., 2020), LLAMA-2 (Touvron et al., 2023), and more.
-Recently, OpenAI created a mechanism (OpenAI, [n.d.]b) to allow users to create custom version of ChatGPT.
-To support this proliferation, sharing, and reuse (finetuning and editing) of large models, many models are hosted on platforms to support the collaborative use and sharing of models such as HuggingFace (Face, 2023) and Kaggle (Kaggle, 2023).
-In the data management community, we have done a lot of introspection about how machine learning and Large Language models will revolutionize data management (Halevy et al., 2023; Fernandez et al., 2023).
-It is true that for many important (and stubbornly hard) semantic data management problems such as data discovery and data wrangling, advanced machine learning models are out performing other approaches (Fan et al., 2023b, a; Dong et al., 2023; Li et al., 2020; Narayan et al., 2022).
-Exciting research challenges remain in applying these models in data management and some point out the lack of exploration of models beyond LLMs (Borisov et al., 2022).
-However, we do not believe this makes machine learning an existential threat to data management or a challenge to the the raison d’être of data management (Fernandez et al., 2023).
-Rather, we argue that the large model revolution poses some new and fascinating scientific challenges that data management researchers are, perhaps, uniquely qualified to tackle.
-In this vision paper, we pose the question of how can principled data management revolutionize the use and management of models, including important problems like model selection (see Example 1) and model provenance.
-While model management (Bernstein et al., 2000) has long been a core topic in data management, we have been slow to move beyond the management of schemas and mappings (Doan et al., 2012) to the management of large machine learning models.
-Today, the way in which we search for models is through 20th century keyword search over (usually manually specified) names or metadata.
-The way in which we understand the relationship between two models is quite primitive (for example, is one manually declared to be a fine-tuned version of another).
-And the study of model provenance or lineage is just emerging (Mei et al., 2022; Mu et al., 2023).
-We draw on lessons learned from collaborative data sharing platforms, such as Data Hub (Bhardwaj et al., 2015), that led to the development of data lakes (Nargesian et al., 2019) as powerful repositories for the management of heterogeneous collections of data.
-We introduce model lakes, as a parallel to data lakes, and discuss how important innovations in data lakes, including data discovery, annotation, and version management, can (and should) be applied within model lakes and studied with the same vigor.
-Example 1 ().
-Model Selection Problem: Consider a situation where a user wants to find a model that can summarize a legal document and simplify it in a non-technical manner.
-On HuggingFace (as of Jan 2024), the user finds that there are around 489,000 models uploaded and around 1500 of them have the ‘summarization’ task tag.
-While there are filters (trending, most likes, most downloads, model name search, and more), the user finds it hard to choose which model they should actually use.
+With the dramatic rise in AI capabilities across a variety of domains (Deng et al., 2023; Schick et al., 2024; OpenAI, [n.d.]; Google, [n.d.]a; AI, [n.d.]a; Microsoft, [n.d.]; AI, [n.d.]b; Labs, [n.d.]b; Google, [n.d.]b; Trummer, 2022; Chintagunta et al., 2021), many organizations have begun to commit significant resources to developing Machine Learning Models.
+Many of these are fine-tuned versions of popular foundation models, such as Llama-3 (Wang et al., 2024b), Mistral (Jiang et al., 2023), DeepSeek-R1 (DeepSeek-AI et al., 2025), Stable Diffusion (Rombach et al., 2022), BART (Lewis et al., 2020), and BERT (Devlin et al., 2019).
+Proprietary closed-source models such as GPT-4 (OpenAI, 2023), Gemini (Anil et al., 2023) and Claude-3 (Anthropic, 2024) also support creation of fine-tuned models.
+To support this proliferation, sharing, and reuse of large models, many models are hosted on platforms to support the collaborative use and sharing of models such as Hugging Face (Face, 2023) and Kaggle (Kaggle, 2023).
+As the number of pre-trained models grows, comparing them and selecting the right one for specific tasks becomes increasingly difficult (see Example 1.1).
+Documentation, particularly model cards (Mitchell et al., 2019), aims to provide essential insights, but Liang et al.
+(2024) have revealed a concerning lack of transparency and completeness in these resources.
+This makes it hard for users to make informed decisions, especially when navigating model sharing platforms.
+Efforts such as the Data Provenance Initiative (Longpre et al., 2023), MLCommons (Reddi et al., 2019; Mattson et al., 2019), and Responsible Foundation Model Development Cheatsheet (Longpre et al., 2024) have been introduced to enhance the documentation of model creation and capabilities, with a particular focus on detailing their training data.
+However, not all model creators adhere to these guidelines, meaning that many existing and future models may still lack crucial information needed by users.
+To address this gap, we propose a systematic breakdown and formalization of tasks for “model lakes" — a system containing numerous heterogeneous pre-trained models and related data in their natural formats (one example is Hugging Face (Face, 2023)).
+We introduce model lakes, as a parallel to data lakes, and discuss how important innovations in data lakes (Nargesian et al., 2019), including data discovery, annotation, and version management, can (and should) be applied within model lakes and studied with the same vigor.
+We look at how model lakes are currently managed and define and discuss tasks that can be used to better inform users about the models and their relationships.
+Others have discussed how one type of model (LLMs) may disrupt data management (Fernandez et al., 2023).
+We focus on how data management can transform the management and use of AI models.
+Example 1.1.
+Model Search Problem: Consider a situation where a user wants to find a model that can summarize a legal document and simplify it in a non-technical manner.
+On Hugging Face (as of Sept 2024), the user finds that there are around 1M+ models uploaded and 1950 of them have the ‘summarization’ task tag.
+While there are filters (trending, most likes, most downloads, model name search, and more), the user finds it hard to choose which model to use.
 There are various concerns that the user goes through as she scrolls through different model cards (a common semi-structured form of model documentation).
@@
 Are they also trained on the same or different legal texts?
-2. Models
-Consider a model, , which is trained with dataset, .
-This model has an algorithm (function), , where .
-Based on the training objective provided by , can be discriminative or generative in nature.
-In the case of discriminative models, would learn to differentiate and classify different kinds of data instances.
-In the generative case, can be a sampling process that produces a distribution of new data instances.
-For instance, text-conditional diffusion models (Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022) take a text input, and produce an output that contain pixel values that form an image that matches the text description.
-For large language models (Touvron et al., 2023; Zheng et al., 2023; Brown et al., 2020), the output can be generated text (which may be in the format of tabular or semi-structured data), that the model creates in response to an input prompt .
-In general, training data can be of various types or combinations — it can be text documents, images, audio files, etc., that may be labeled or unlabeled.
-The full algorithm would include both ’s architecture and a pipeline for training and testing it.
-could be a parameterized model with any architecture, such as a Convolutional Neural Network (Cun et al., 1990) or a Transformer (Vaswani et al., 2017) or a simpler model.
-Note that unlike others (Fernandez et al., 2023), in our work, we consider not just LLMs, but also the worker-bee models that may be discriminative or generative, and that have been driving data science innovation since before the ChatGPT era.
-Moreover, recent findings from a survey on deep learning and tabular data (Borisov et al., 2022) underscore the necessity of investigating novel architectures for adoption, including diffusion or hybrid classical machine learning and deep learning models.
-They stress this is particularly important for handling heterogeneous collections of tabular data with better predictive performance and computational complexity.
-By developing and advancing a semantic model management system, the exploration of these models can be conducted in a more informed and efficient manner.
-3. State-of-the-art in Model Management
-Model Repositories and Model Registries.
-The proliferation of models has lead to the emergence of model repositories such as Civitai (Inc., 2023), HuggingFace (Face, 2023), or Kaggle (Kaggle, 2023) that openly share models.
-Companies, such as AWS also provide model repository products (such as the Amazon Elastic Container Registry) that can be used to share models within an organization.
-These platforms let users search for models by browsing through catalogs or by doing keyword search over manually created model names (and sometimes metadata).
-Beyond the hosting and sharing of models, a model registry provides clear naming conventions (for models and versions) and standardized, structured metadata (examples include Neptune (Labs, [n.d.]a) and AWS’ Sagemaker Model Registry (Verma et al., 2022)) and is more of a comprehensive system that enables documentation of the entire lifespan of machine learning models.
+2. Three Viewpoints of a Model
+An AI model, , can be analyzed from three viewpoints: according to its history, intrinsic composition, or extrinsic behavior.
+These viewpoints highlight different aspects of the model’s characteristics, aspect that we will show are useful in analyzing model lake problems and solutions.
+The history of the model is defined by its training data () and training algorithm (), which may include processes like fine-tuning, model editing, or other adaptation techniques.
+The degree to which history is documented within a model lake can vary greatly (Liang et al., 2024).
+The intrinsic viewpoint concerns the model’s internal structure.
+This includes the model architecture (), and the specific trained parameters ().
+The architecture refers to the structural design of the model, such as a combination of multi-layer perceptrons (MLPs) and attention layers in transformer models.
+Formally, the architecture defines a function family which is instantiated with the specific learned parameters to create the function
+In contrast, the extrinsic viewpoint focuses on the model’s observable behavior and performance in user-defined tasks.
+The model parameters, , are not visible as part of the extrinsics.
+However, the function and the model behavior, , is extrinsic.
+For example, in the case of an unconditional generative model, the extrinsics correspond to the observable probability distribution defined by the model, .
+For a classifier, the extrinsics are defined by the behavior of the modeled predictions .
+Extrinsics can be observed in terms of a neural network’s action on inputs and outputs and , without requiring any knowledge of its training data or its internal structure.
+The distinction among intrinsics, extrinsics, and history is useful because, while every model has all these characteristics, there are cases where certain aspects may be unavailable.
+For example, in a Model Lake, the intrinsic details of some models might be inaccessible, and some analysis methods may rely solely on extrinsic observations or historical records to understand a model’s behavior.
+We use this distinction to analyze possible solutions to a variety of model lake tasks.
+Our envisioned model lake is depicted in Figure 1.
+3. Formalizing Model Lake Tasks
+Model lake tasks are specifically concerned with gathering and presenting insights about the models that are stored within the lake.
+Topics related to the infrastructure behind model training or storage systems fall outside this scope.
+Model Attribution.
+In data lakes, data provenance is the “description of the origins of a piece of data and the process by which it arrived in a database" (Buneman et al., 2001).
+Model provenance (often called attribution) considers questions like “Why was image X generated when the model was given input Y?" If the history is recorded, the history can be consulted to directly ask the training data attribution question formally: which training data items are most influential on the decision; in other words, which , if they were not present in the training data, would cause the decision to change the most?
+When history is not available, attribution can be studied using intrinsic and/or extrinsic clues to provide insight into the attribution of model decision behavior.
+For example, we can perform sensitivity analysis on the observable extrinsics of a model by asking: which aspects of the inputs to or are most important in a model’s prediction of a particular output?
+And if we have access to model intrinsics, we can study feature or representation analysis, which asks, which internal representations or internal “concepts” within the model are most important for a decision?
+Model Versioning.
+In the model versioning task, we want to understand whether (and possibly how) a model has been created from other models (similar to studying how a version of a data set or table may have been derived from another (Shraga and Miller, 2023)).
+One possible definition of model versioning that uses an intrinsic viewpoint (Horwitz et al., 2024) is: Given a model, and a set of models, , construct a directed Model Graph, , where a directed edge between models indicates that one model is a version of the other.
+The edges can describe the transformation.
+This can include training techniques, optimization techniques, as well as the data used to further train (fine-tune) models.
+An important problem in versioning is: given a model’s training parameters and a candidate model’s parameters, , is a source of ?
+Being a source model would mean that or a version of was used for training .
+Model Search.
+Model search is the task of finding a related or desired model.
+Again we can consider this task from different viewpoints.
+An extrinsic view considers the behavior of the model.
+Given a task function, where a task takes an input, and produces an output , we want to find the best performing model, , for the given task among a set of candidate models, .
+Each candidate model is characterized by their observable behavior, i.e., .
+The optimal model is selected based on a scoring function of how close the model’s behavior is to a query model.
+Even considering only an extrinsic view, there are many formulations of model search.
+If the goal is to find models that perform similarly on specific data (for example, a specific image), then a possible definition is (Lu et al., 2023): given a point, , and a set of candidate models, , rank models based on their similarity of their respective observable behavior w.r.t.
+to .
+This could be extended to other types of models where for a given model, and a set of candidate models, , we want to find related models based on a ranking function that considers the similarity of models’ output distribution and semantic concepts/patterns present within them.
+An intrinsic view of model search would find models with similar model architectures and training parameters.
+Or as done in data lakes where intrinisic search is the norm, we could create embeddings representing the important features of the model and design a fast nearest neighbor search over these embeddings.
+Considering model history, we could search for models that have been trained on the same dataset.
+This is straight-forward when history is recorded, but when it is not fully explicit, we may leverage extrinsic or intrinsic clues in the search.
+Benchmarking.
+For single model tasks, a benchmark, (for example, a set of images), is used to measure the performance of a model, (or set of models) based on a scoring function, .
+For model lake tasks, we will need new (shared) model lake benchmarks (large sets of models ) that can be used to measure the quality of a lake task solution.
+This means that within a benchmark lake, we will need verified ground truth.
+4. State-of-the-art in Model Mgmt
+Model Repositories, Registries, and Lakes.
+A model repository is a storage system for machine learning models.
+A model registry goes beyond basic storage by offering version control (the representation of versions, not today the discovery of version relationships mentioned in Section 4).
+These registries typically enforce clear naming conventions (for models and versions) and organize structured, standardized metadata (Labs, [n.d.]a; Verma et al., 2022).
+Recent advancements have added functionalities like explainable network intrusion detection (Spadari et al., 2024) for security and integrations of different tools such as model monitoring and experiment tracking into a single framework (Vishwambari and Agrawal, 2023).
+Model registries are typically private, facilitating collaboration within organizations, while open platforms such as Civitai (Inc., 2023), Hugging Face (Face, 2023), and Kaggle (Kaggle, 2023) enable public sharing of models.
+These platforms, which we can consider as model lakes, assist users in model exploration by providing curated catalogs and keyword search, often leveraging both manually created model names and metadata for more efficient discovery.
+Unlike lifecycle creation and training platforms, model lakes focus on managing a set of AI models, including their interrelationships.
+This is distinct from model management in databases, which deals with schemas and their mappings (Bernstein and Melnik, 2007).
+Current model lakes capabilities lack effective mechanisms for representing and navigating the model space semantically, particularly when model documentations or names are incomplete or unknown to users, leaving many models effectively undiscoverable (Liang et al., 2024).
 Documenting Models.
 Models that are deployed are usually accompanied with documentation known as model cards (Mitchell et al., 2019).
-This report generally contains the following categories of information.
-Model Details: Information about the model such as model date, version, type, training algorithms used and other resources to learn more about the model.
-Intended Use: Use cases that were imagined while developing the model and related primary users.
-Factors: Summary of model performance across relevant features (like gender in people-centric models), instrumentation (like recording instrument for audio-based models), and environment (like image lighting for image-based models).
-Metrics: Details of model performance measures, decision thresholds, and approaches to uncertainty and variability.
-Evaluation Data: Information on dataset(s) used for quantitative analyses, the motivation behind choosing the dataset(s) and related preprocessing methods.
-Training Data: Similar to information present for evaluation data section, but on data used to train the models.
-Quantitative Analyses: Results of evaluating a model based on the metrics, which may be broken-down based on factors mentioned earlier.
-This includes unitary, i.e., model performance with respect to each factor, and intersectional results, i.e., model performance with respect to the intersection of factors.
-Ethical Considerations: Information on ethical concerns such as model usage of sensitive data, intention of using the model for human life-related decisions, potential ethical risks and harms.
-Caveats and Recommendations: Details of concerns or suggestions for using the model that are not covered elsewhere.
+Model cards contain (among other categories) information on model details, intended use, metrics, training data, and quantitative analyses.
 Similar to datasheets (Gebru et al., 2021), model cards are designed to guide developers in documenting models in a structured way.
 Model cards can be and should be augmented with information more similar to nutritional labels (Stoyanovich and Howe, 2019) that also include information about fairness and bias in the data (models).
-They can be further enriched with lineage and security related documentation such as adversarial attack and related defense measure policies, which are outlined in FactSheets for AI services (Arnold et al., 2019).
+They can be further enriched with lineage and security related documentation (Golpayegani et al., 2024) such as adversarial attack and related defense measure policies, which are outlined in FactSheets for AI services (Arnold et al., 2019).In addition to research on enhancing the completeness of model documentation, there remains a critical gap in the verification of model cards.
+There is a danger that people could intentionally misinform model users with malicious intent (Security, [n.d.]b).
+The state-of-the-art in verifying the documentation of a model is notably in its infancy (Longpre et al., 2024).
+Mithril AICert (Security, [n.d.]a) is developing a certification initiative that verifies whether a model was trained using the specified algorithms and data sets.
+However, this initiative has some limitations.
+First, as it is still in development, it is not yet available for production use.
+Second, the certification process depends on the voluntary participation of the model creators.
+Furthermore, the AICert website (Security, [n.d.]c) highlights several critical limitations, including the inability to audit training code or data for risks such as backdoors or data poisoning.
 Model Search and Discovery.
-In Example 1, we showcase a potential scenario for the Model Selection Problem.
-The current solution pipeline involves a user searching for a relevant model by naming specific models or by typing related keyworks like legal to find models that either have that word in their name or in their model card.
+In Example 1.1, we present a potential scenario for the Model Search Problem.
+The current solution pipeline involves a user searching for a relevant model by naming specific models or by typing related keywords like legal to find models that either have that word in their name or in their model card.
 In other words, the search relies on the model’s name and documentation.
-Hence, any sorting of the answer by relevance is just the relevance (prevalence) of the keywords and is not a semantic notion.
+Hence, any sorting of the answer by relevance is just the relevance (prevalence) of the keywords and is not a semantic notion based on the model itself.
 Of course, this search may fail if the documentation is incomplete or inaccurate.
-The state-of-the-art in verifying the documentation of a model is also in its infancy.
-Interpretability.
-Interpretability has been defined as the extraction of relevant knowledge from a model concerning relationships either contained in data or learned by a model (Murdoch et al., 2019).
-Interpretability is critical for model analytics to enable semantic analysis and understanding of models and their training data.
-Several works have surveyed techniques and research questions in this area (Singh et al., 2024; Luo and Specia, 2024; Zhao et al., 2024; Räuker et al., 2022; Doshi-Velez and Kim, 2017; Gilpin et al., 2018), which can be broadly categorized into the following areas: Local model explanations explain the sensitivity of individual output predictions to local changes in inputs using gradients (Sundararajan et al., 2017), masks (Petsiuk et al., 2018; Fong et al., 2019), local models (Ribeiro et al., 2016), or Shapley values (Lundberg and Lee, 2017; Chen et al., 2023).
-Global explanations explain the mechanisms of the overall model at the level of attention patterns (Clark et al., 2019), representations (Belinkov, 2022; Zou et al., 2023), circuits (Elhage et al., 2021; Wang et al., 2022), or neurons (Bau et al., 2018, 2017; Geva et al., 2023).
-Models can be designed to be inherently interpretable (Rudin et al., 2022).
-Lastly, datasets can be explained using natural language explanations (Narayan et al., 2022), data visualizations (Dibia, 2023), or by training inherently interpretable models (Ustun and Rudin, 2016; Singh et al., 2023) of the data set.
-Provenance.
-A new and important line of inquiry considers questions related to provenance (Mei et al., 2022; Mu et al., 2023).
+Within open data and enterprise data, researchers have learned that they cannot rely on metadata for datasets to be accurate, complete, or consistent within a data lake (Nargesian et al., 2019).
+As a result, there is a great deal of work on semantic or content-based search in data lakes including join search (Zhu et al., 2016, 2019; Khatiwada et al., 2022; Dong et al., 2023), union search (Nargesian et al., 2018; Khatiwada et al., 2023; Fan et al., 2023b; Hu et al., 2023), and related dataset search (Fernandez et al., 2018; Bogatu et al., 2020).
+Machine learning models have revolutionized content-based dataset search.
+Important and impactful work has shown how we can use machine learning to create meaningful semantic representations of tuples, columns, and full tables to enhance dataset search and other semantic tasks like data integration and alignment.
+But what about content-based model search?
+To the best of our knowledge, this is an area within ML that is in its infancy.
+Recent techniques for image models leverage meta learning (Lu et al., 2023; Li et al., 2023).
+HuggingGPT (Shen et al., 2024) uses an LLM (in their case, ChatGPT) as a controller to decide which models to use based on a user’s prompt.
+This metadata-based search differs from content-based approaches, as the LLM parses the user’s prompt into tasks and uses model descriptions to select relevant models.
+While it allows queries across any modality or domain, it is limited by the LLM’s capabilities and the quality of model documentation.
+Additionally, it may fail when object-centric tags do not fully capture the model.
+Our vision emphasizes that content-based model search must cover all models in model lakes, including large language models, while ensuring usability through speed and accuracy.
+Another important problem is related model search.
+One approach to addressing related model search was explored by Lu et al.
+(2023), who search for image generative models by using the behavior of another image model as a query.
+We propose extending such model as query searches to the model lake to help users identify related models when exploring the model in question.
+Attribution.
+An important line of inquiry considers questions related to provenance or attribution (Mei et al., 2022; Mu et al., 2023).
 These issues pose similar questions to those studied in the database community (Buneman and Tan, 2018): from where did a generated fact derive or why was a predication made?
-A subset of these questions are also posed in model interpretability as part of local and global explanations.
-As in data provenance, the main issue is in whittling down all the provenance associated with some process (such as a model) into simple, but useful explanations (Buneman and Tan, 2018).
+Building on the concept of data provenance (Buneman et al., 2001), we can extend the notions of why-, where-, and how-provenance to models, under the name of model attribution.
+Similar research questions are also posed in model interpretability as part of local and global explanations.
+As in data provenance, the main issue is in whittling down all the provenance associated with some process (such as a full model) into simple, but useful explanations (Buneman and Tan, 2018).
 More than two decades of research in the data management community have produced elegant and simple, yet powerful, models for data provenance when the process is a query (Buneman et al., 2008, 2016) or workflow (Moreau et al., 2013; Moreau and Groth, 2013).
-In large model provenance, the goal cannot be to understand and track all inputs (data, hyperparameters, code, training regimes, …) that were used to create a specific model output.
-Rather research is needed on finding meaningful sets of concepts that can be tracked efficiently and which provide important insights into model behavior.
-There has been some initial research on solving the model provenance.
-Mu et al (Mu et al., 2023), for instance, uses a data-driven and model-driven representation learning method to encode ‘Model DNA’, which the authors then use as part of a framework for model provenance identification.
-Their goal is identify whether a model is a pre-trained model of another model.
-The former model is known as the source model while the latter model is known as the target model.
-This technique assumes a relaxed condition where the source and target model have the same architecture or structure and the target model’s training is based entirely on the source model.
-Furthermore, it assumes that we can access training data of the source model.
-However, there are harder conditions that can exist where the target model is trained based on a subset of the parameters of the source model or when the architecture of the target model is different from that of the source model.
-We define the granularity of such model conditions in Section 5 and explore various ways a model can be different from one another.
+In large model attribution, the goal cannot be to understand and track all inputs (data, hyperparameters, code, training regimes, …) that were used to create a specific model output.
+Hence, the challenge is to find meaningful sets of concepts that can be tracked efficiently and that provide important insights into model behavior.
+The training data attribution problem is nontrivial because every aspect of training has the potential to impact every decision of a model.
+A variety of methods have been developed to estimate influence of training data on model behavior (Koh and Liang, 2017; Grosse et al., 2023; Wang et al., 2023).
+The current approaches require extensive use of training data as well as costly analysis of model intrinsics.
+Another approach to the problem is to apply techniques such as membership inference analyses, or membership inference attacks (Shi et al., 2024) that ask the question of whether a specific training data item is present in the training data , or training data reconstruction methods that extract sets of items from the training data (Carlini et al., 2021; Schwarzschild et al., 2024).
+Insight about the attribution of model behavior can also be studied through model interpretability methods: extraction of relevant knowledge from a model concerning relationships either contained in data or learned by a model (Murdoch et al., 2019).
+Several works have surveyed techniques and research questions in this area (Mueller et al., 2024; Singh et al., 2024; Luo and Specia, 2024; Zhao et al., 2024; Räuker et al., 2022; Doshi-Velez and Kim, 2017; Gilpin et al., 2018).
+These approaches can be broadly categorized into the following areas.
+Local model explanations explain the sensitivity of individual output predictions to local changes in inputs using gradients (Sundararajan et al., 2017), masks (Petsiuk et al., 2018; Fong et al., 2019), local models (Ribeiro et al., 2016), or Shapley values (Lundberg and Lee, 2017; Chen et al., 2023).
+Global explanations explain the mechanisms of the overall model at the level of attention patterns (Clark et al., 2019), representations (Belinkov, 2022; Zou et al., 2023a), circuits (Elhage et al., 2021; Wang et al., 2022), or neurons (Bau et al., 2018, 2017; Geva et al., 2023).
+Models can also be designed to be inherently interpretable (Rudin et al., 2022).
+Lastly, datasets can be explained using natural language explanations (Narayan et al., 2022), data visualizations (Dibia, 2023), or by training inherently interpretable models (Ustun and Rudin, 2016; Singh et al., 2023).
 Model Versions.
-Models are valuable assets that are reused in numerous ways to create new versions.
-Based on the training type, the function or algorithm, can vary.
-For instance, if a model is pre-trained, would include a pipeline were it would encourage to learn general features from dataset .
-However, if is fine-tuned, would start off with weights from a previously trained model and learn features from to understand a specific task better.
-While pre-training and fine-tuning are some of the most common training objectives for creating new versions of a model, there are newer, more surgical-like training objectives such as model editing (Sinitsin et al., 2020; Meng et al., 2022) where focuses on updating certain facts (for instance, updating the name of the current President of a country) learned by a model in a more localized fashion.
-With various options for training methods, architecture, and dataset combinations, enormous numbers of models have been created with varying degrees of similarities and differences amongst each other.
-Among the versioning methods, we examine fine-tuning and editing in more detail.
-Fine-tuning is the process of further training the parameters of an existing model to improve its performance, usually on specific task(s) or domain(s).
+Models are valuable assets that can be adapted and reused to create new versions.
+The original or base version is typically a foundation model — a pre-trained model that learns general features from its training dataset, denoted as .
+By making further adjustments to the training algorithm (), architecture (), or dataset (), subsequent versions of the base model can be developed.
+Common -based modifications include fine-tuning, parameter-efficient tuning, preference tuning, model stitching, and model editing.
+Model stitching, for example, involves altering by combining the architectures of two or more models to create a hybrid model (Lenc and Vedaldi, 2015).
+Model editing (Gandikota et al., 2024; Orgad et al., 2023; Meng et al., 2023, 2022; Mitchell et al., 2022; Sinitsin et al., 2020) focuses on updating certain facts (e.g., changing the name of the current President of a country), making localized adjustments without retraining the entire model.
+Fine-tuning involves further training a model’s parameters,, to improve performance on specific task(s) or domain(s).
 For instance, T5 (Raffel et al., 2020) is a pre-trained model that has been trained on a collection of large collection of text (about 750 GB) to perform well on a diverse set of tasks.
 This has been further fine-tuned with the MIMIC-III (Johnson et al., 2016) and IV (Johnson et al., 2023) dataset to form Clinical-T5 (Lu et al., 2022) to perform better for medical domain-related tasks.
-Another model instance is FLAN-T5 (Chung et al., 2022) that has been fine-tuned on data with 1.8K tasks phrased as instruction so that it can handle such prompts better.
-Apart from datasets and training goals, there are also various optimizations schemes that could differentiate various fine-tuned models.
-For instance, Hu et al.
-(2022) experimented with a GPT-2 model fine-tuned with Adam (Kingma and Ba, 2015) and another GPT-2 model instance fine-tuned with their Low Rank Adaption (LoRA) (Hu et al., 2022).
-They both have the same task goals and are fine-tuned from the same source model, but have a difference in performance due to the fine-tuning optimization conducted.
-Consequently, with different datasets, training goals and algorithms, various versions of a pre-trained model can be created.
-Because they have the same source model, the fine-tuned models can also be considered to be related to each other.
-Direct model Editing.
-Large language models (LLMs) contain a wealth of knowledge encoded in their parameters.
-As a result, they can respond to natural-language queries such as those that relate to factual statements.
-However, LLMs can sometimes hallucinate or point to outdated information.
-Hence, there have been several studies that look into localizing and understanding the internal mechanisms within LLMs.
-Zhu et al.
-(2020), for instance, proposed the task of modifying factual knowledge in transformer models such as BERT (Devlin et al., 2019) by performing a constrained fine-tuning on the original model with specific facts.
-De Cao et al.
-(2021), on the other hand, introduced Knowledge Editor, that uses hyper-networks (Ha et al., 2017) to tune the model’s parameters for editing specific knowledge.
-In terms of finding these facts, Dai et al.
-(2022) proposed the idea of knowledge neurons, whose activation highly correlated with their related knowledge expression.
-MEND (Mitchell et al., 2022) also edits local knowledge, but it does so by training a collection of Multi-Layer Perceptrons (MLP) to modify model gradients that would result in the production of local model edits.
-In contrast, ROME (Meng et al., 2022) uses rank one modification of the MLP weights of a single layer to directly edit appropriate weights that would change the fact associated to an information present the transformer model.
-MEMIT (Meng et al., 2023) builds upon ROME to be able to edit multiple knowledge expressions in the model.
-Direct model editing can also be applied to multimodal models: the TIME (Orgad et al., 2023) and UCE (Gandikota et al., 2024) methods edit text-to-image diffusion models.
-All these methods showcase various ways to edit a model, and hence, can be useful to understand how a current model is changed from its base model when a fact is changed.
-4. Model Lakes: Vision
-Given this state-of-the-art, we now outline our vision for model lakes: powerful repositories for the management of heterogeneous models.
-We enumerate challenges that require new scientific solutions and provide evidence that these challenges can be met.
-Content-based Model Search.
-In analogy with dataset search, Google provides a comparable service to model registries, but for datasets (Benjelloun et al., 2020; Alrashed et al., 2021).
-They require dataset owners to follow their naming conventions and use metadata from Schema.org.
-Note this type of search only works well when there are clear incentives for artifact owners to provide complete and compliant metadata.
-Google provides this by returning datasets that comply with its best practices, though even Google struggles with impostors who game the system (Alrashed et al., 2021) by giving invalid semantic annotations.
-Current model repositories, such as Hugging Face, have found that they are subject to model designers who give sparse or ambiguous metadata within their model names or model cards (Liang et al., 2024).
-There is a danger that people could intentionally misinform with malicious intent (Security, [n.d.]b).
-Within open data and enterprise data, researchers have learned that they cannot rely on metadata for datasets to be accurate, complete, or consistent within a data lake (Nargesian et al., 2019).
-As a result, there is a great deal of work on content-based search in data lakes including join search (Zhu et al., 2016, 2019; Khatiwada et al., 2022), union search (Nargesian et al., 2018; Khatiwada et al., 2023; Fan et al., 2023b), and related dataset search (Fernandez et al., 2018; Chapman et al., 2020).
-As mentioned in the introduction, machine learning models have revolutionized content-based dataset search.
-Important and impactful work has shown how we can use machine learning to create meaningful semantic representations of tuples, columns, and full tables to enhance dataset search and other semantic tasks like data integration and alignment.
-But what about content-based model search?
-There is a bit of “physician, heal thyself” here.
-To the best of our knowledge, this is an area in its infancy.
-Lu et al.
-(2023) are the first to introduce the task of content-based model search, ”which aims to find the most relevant deep image generative models that satisfy a user’s input query”.
-Their solution builds a representation of the output distribution of a generative image model and uses this to find models that are most relevant to an input image or to an input model (or more precisely to its distribution).
-The trend of using meta-learning (and creating a hyper-network) for this task continues in other methods, such as AutoMRM, which is a model retrieval method based on multi-modal query and meta-learning (Li et al., 2023).
-A current limitation of these methods are that they are only able to perform retrieval for models that generate images.
-HuggingGPT (Shen et al., 2024) overcomes this limitation by using an LLM (in their case, ChatGPT) as a controller to decide which models to use based on a user’s prompt.
-This, however, is a metadata-based search and not a content based one, since the LLM would parse the user’s prompt into solvable tasks and then utilize another input of model descriptions (metadata) to decide which models would be relevant for each task.
-This approach allows queries and solutions in any modality or in any domain.
-However, it is limited by the capabilities of the LLM employed and the quality and accuracy of the model documentation.
-Furthermore, unlike content-based search, metadata-based ones would primarily fail when object-centric tags do not describe the model fully.
-As part of our vision, it is crucial that the content-based model search methods cover all types of models in the model lakes, including large language models.
-Additionally, ensuring their commercial usability, driven mainly by speed and accuracy, is essential.
-A hybrid approach combining metadata and content-based search could be effective, especially when exhaustive and verified metadata is accessible.
-Related Model Search.
-Another important problem is related model search.
-A model is considered similar or related to another if they have some degree of commonality in one or more combinations of the following categories: datasets, training algorithms, behavior, or purpose.
-One approach to addressing related-model search was explored by Lu et al.
-(2023), which searches for image generative models by using the behavior of another image model as a query.
-We propose extending such model-query searches to all models in a model lake to help users identify related models when exploring the model in question.
-This can then be used as an informative analytics visualization as shown in Figure 1.
-Documentation Verification and Auditing.
-It is crucial to be able to verify that manually created documentation is correct and complete.
-When documentation is missing or incomplete, can we learn it?
-This is similar to discovering and annotating metadata in data lakes (Sawadogo and Darmont, 2021; Zhao et al., 2021; Korini and Bizer, 2023; Fan et al., 2023a).
-If there is a documentation, the goal would be to verify them and add any information that was left out.
-Currently, Mithril AICert (Security, [n.d.]a), for instance, runs an initiative to provide certification to indicate that a model is trained with the specified algorithm and with the specified training dataset.
-However, there are a couple of caveats.
-First, this initiative is still under development and hence, cannot be used for production at the time of writing.
-Second, this mechanism is dependent on creators signing up to gain such certification.
-Furthermore, based on the limitations provided on AICert (Security, [n.d.]c) website, the training code and data itself cannot be audited for threats such as “backdoors injected into a model by the code or poisonous data.” Lastly, it also states that AICert also needs to be inspected since we are currently trusting that it is running at a safe server and does not have a malignant app from which users have to run their code on the training data.
-Hence, there are a couple of broad research questions to explore within the document-centric verification realm – What information do we need for a model to be reproducible?
-If the information is provided, how can we verify the information without being compromised in the process?
-If the information is not provided, how can we gather the necessary information from the models themselves?
-How can we automate this process?
-While developing models, there are systems like MISTIQUE (Vartak et al., 2018), that allows developers to perform model diagnosis, i.e., the process of examining machine learning model performance, and aiming to understand where the model excels and where it falls short.
-MISTIQUE can do so by capturing and storing the model intermediates, which are generally data artifacts that relate to the model including input data, prediction values, and data representations produced by model.
-These intermediate values can then be queried by the developers to diagnose the model to understand its performance.
-However, such frameworks are limited to white-box models, which are models where users can have full access including access to weights, activation, and gradients.
-However, we also need to consider documentation verification for other variations of model access such as black-box access, where users can only see the model’s output for their queries.
-Currently, there is no rigorous scheme present, especially for AI auditing (Casper et al., 2024).
-Data Citation.
+In addition to the traditional fine-tuning strategy, parameter-efficient fine-tuning methods have emerged to reduce computational overhead by freezing most of the model’s parameters, only updating a small subset necessary for fine-tuning.
+For instance, Low Rank Adaption (LoRA) (Hu et al., 2022) is a parameter-efficient fine-tuning method to adapt various models by only updating a low-rank subset (Hu et al., 2022).
+Preference tuning is another advanced technique, as seen in ChatGPT (OpenAI, [n.d.]) and InstructGPT (Ouyang et al., 2022), which integrates human feedback into the fine-tuning process (Winata et al., 2024).
+Lastly, due to the emergent ability of models to perform tasks without training (Dong et al., 2022; Liu et al., 2023), newer models leverage prompting as a way to control content generation without needing further updates to their parameters.
+Information about model versioning can be inferred even if the model history is unavailable.
+For example, Mu et al.
+(Mu et al., 2023) propose a data- and model-driven method to encode “Model DNA" for identifying if a model is a pre-trained version of another, assuming both share the same architecture and training data.
+However, more challenging cases arise when the target model has a different architecture or is trained on only a subset of the source model’s parameters.
+Hugging Face recently introduced new metadata fields in their model cards, enabling users to specify the base model and explain how it has been modified.
+This metadata generates a model tree, linking related models by their extensions.
+However, its accuracy depends on reliable documentation, and older models lack this data.
+Research on reconstructing and verifying relationship is emerging (Horwitz et al., 2024; Sun et al., 2025; Zhu et al., 2025), such as Horwitz et al.
+(2024)’s approach using weight similarities, though this approach is limited to known models with a single base version.
+Privacy and Safety.
+Models are vulnerable to the disclosure of private information (Shokri et al., 2017; Carlini et al., 2021) and adversarial results (Zou et al., 2023b; Li et al., 2025) when attacked.
+Hence, initiatives such as Privacy Preserving Machine Learning (Ruehle et al., 2021) exist to understand, measure, and mitigate such risks.
+As a result, techniques such as differential privacy (Dwork et al., 2006; Yan et al., 2024), data sanitization (Staab et al., 2025; Dou et al., 2024) and robust prompt optimization (Zhou et al., 2024) have been utilized to defend against such attacks.
+These methods generally aim to obscure or eliminate private information while detecting and preventing attempts to jailbreak or manipulate the model and its output.
+However, this can create a false sense of privacy as defense techniques can continue to be compromised with other attack schemes (Xin et al., 2024).
+Beyond privacy, ensuring AI systems are fundamentally safe is another critical challenge.
+Community-driven efforts focus on building safe AI (Ortega et al., 2018), aiming to specify, verify, and ensure that the models behave as intended.
+A recent approach leverages neuroscience concepts, such as representation engineering, to enhance AI transparency (Andy Zou, 2023; Mineault et al., 2024) and improve our understanding of traits like honesty, power seeking tendencies, and morality in models.
+However, this direction is still in its early stages, with open challenges including determining the range of representations of concepts and functions that can be extracted from models and developing scalable evaluation methods to better inform users about model safety.
+Benchmarking.
+Benchmarking plays a crucial role in evaluating model performance for specific tasks and remains a well-established research area, however benchmarks for newer topics such as model attribution and versioning are urgently needed.
+Developers frequently report model performance using standardized benchmarks, making it a routine part of model assessment.
+Broadly, there are two primary types of benchmarking: (1) evaluating how well a model, , approximates a ground truth distribution, and (2) assessing a model’s performance against a targeted evaluation metric.
+In classification tasks, models are typically evaluated using accuracy, as ground-truth labels provide a direct assessment of the model’s output.
+Additionally, confusion matrices offer insights into the types of errors made.
+For text generation tasks, perplexity is a widely used metric, and several popular benchmarks exist to assess performance (Hendrycks et al., 2021a, b; Rein et al., 2024).
+In the case of image generation, the Fréchet Inception Distance (FID) (Heusel et al., 2017) is a common metric and COCO (Lin et al., 2014) and VQAv2 (Goyal et al., 2017) are some notable benchmarks in this domain.
+Beyond traditional performance metrics, benchmarking also considers biases related to protected attributes (Schramowski et al., 2023), as well as the environmental impact (Lacoste et al., 2019) of model training and deployment.
+Model lake benchmarks lack large-scale, publicly available datasets that mimic realistic conditions of diverse models in model lakes.
+There has been preliminary efforts by Lu et al.
+(2023), where they released a benchmark for model search with 259 publicly available image generative models and 1000 customized text-to-image diffusion models.
+Model Zoo (Schürholt et al., 2022) is another relatively large-scale dataset, but it is limited to vision models.
+Creating large-scale model lake benchmark that integrates various modalities (such as text, images, and audio) remains a research challenge.
+5. Research Roadmap
+Building on current research and identified gaps outlined in Section 4, this section presents a vision to address these challenges.
+A model lake is illustrated in Figure 2, where rather than APIs, we envision data scientists interaction with the lake using queries.
+Benchmarking.
+As discussed in Section 4, benchmarking, while a well-established research area, is under-explored in model lakes.
+An important topic is the development of lifelong benchmarks (Prabhu et al., 2024) that can address increasingly complex and novel scenarios as models continue to evolve in capability and diversity.
+In addition, there is a need to develop benchmarks specific to model lake tasks.
+For instance, to advance research in model attribution, a comprehensive benchmark dataset is needed—one that includes labeled model parameters, architectures, and detailed transformation records (e.g., fine-tuning, model editing).
+This benchmark can be extended to model versioning by adding data that tracks the previous and subsequent versions of models.
+Model Inference.
+Deploying effective solutions for model lake tasks is a challenge with respect to usability and scalability.
+To improve usability, the model inference component involves identifying appropriate benchmarks and generating relevant prompts, as well as selecting suitable models (target models or meta-models) to address a user query.
+While users can manually run prompts and select models, this approach is prone to errors and suboptimal outcomes, especially if users lack the expertise to use models effectively.
+For example, a classifier’s behavior may be misinterpreted if a user does not understand the type of data is was trained or the input it expects.
+To mitigate this, the proposed model lake tasks can incorporate additional perspectives, such as intrinsic model properties (e.g., weights), to provide insights and guide users towards a more accurate interpretations and applications.
+This search and generation process can also be automated using an AI agent.
+By applying benchmarks and model(s) in question, we combine research efforts of benchmarking and attribution questions, which can explain a model’s behavior and output.
+Indexer.
+A central component of a model lake is the indexer, which would be used to embed and provide scalable sublinear search over the model embeddings.
+The indexer can use different ranking functions tailored to the task.
+Wang et al.
+(2024b); Achille et al.
+(2019) have introduced an approach for generating model and task embeddings.
+However, search methods must scale to handle millions of models and adapt to newer models with advanced capabilities (Wei et al., 2022).
+Indices like HNSW (Hierarchical New Small World) (Malkov and Yashunin, 2020), have proven effective in practice in indexing high-dimensional embeddings enabling fast nearest-neighbor search (including over data lakes (Fan et al., 2023b).
+However, HNSW provides no formal guarantees on correctness and its use in model lakes remains under-explored.
+Effective embedding of models is crucial for accurate comparison and ranking by the indexer.
+Wang et al.
+(2024b) explores this but there work is not inclusive of all model types.
+A robust system should support diverse embeddings to ensure indexing effectiveness.
+Many of the model lake tasks will benefit from hybrid approach, that indexes both metadata and model embeddings – for example, related model search can combine well-chosen model embeddings representing important intrinsic model features with search over verified models cards.
+Similarly, in versioning, embeddings and their associated rankings can aid in identifying parent-child relationships and assessing the distinctiveness of each model relative to others.
+Weight-Space Modeling.
+Weight-Space modeling is a hyper-representation learning approach where a neural network is trained to process weights of other models (Eilertsen et al., 2020).
+This method can be useful for making distinctions between models, especially in complex scenarios like model stitching, where similar models with multiple shared parent models need to be distinguished.
+This is a promising direction.
+Zhou et al.
+(2025), for instance, reveals a linear connection between fine-tuned models.
+This approach could also facilitate dynamic selection of benchmarks for performance measurement by learning from previous relationships between datasets and models.
+The primary challenge in weight-space modeling is identifying the most relevant aspects of the model for training another model to uncover patterns, while simultaneously considering the weight-space model architecture (Navon et al., 2023).
+It is also crucial to ensure sufficient data diversity to avoid overfitting or underfitting.
+Interpretability.
+Interpretability methods can be most useful for the model attribution task where users must understand the origin of model behavior, for example, when detecting knowledge changes (Youssef et al., 2024), or when unlearning learned knowledge (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2023; Belrose et al., 2023; Kumari et al., 2023).
+Attribution questions also apply to the source model’s architecture.
+Approaches like circuit discovery (Wang et al., 2022; Conmy et al., 2023; Elhage et al., 2021) can help identify the computational origin of model behavior.
+Model inversion can be used to recover an input prompt given an output (Morris et al., 2023, 2024; Huang et al., 2024).
+These methods are also a promising route for understanding the impact of training data on internal model states.
+In recent work, Sharkey et al.
+(2025) present a list of open questions in interpretability, many of which are relevant to the problem of model lakes.
+Holistic Management of Models and Data.
+Effective model lakes need to integrate advances in data lakes in a holistic way given the reliance of models on data as their fuel.
+Many model management tasks include (in part) an analysis of training or input data.
+Hence, traditional data lake concerns such as data provenance, data versioning, and related issues still apply.
+Thus, in addition to new tasks specific to model lakes, we must also account for data lake tasks when dealing with the data used by or generated from these models.
+And integrating these tasks will be important.
+As a simple example, when searching for models trained on a dataset, users may want to find models trained on versions of the dataset.
+6. MODEL LAKE TASKS: APPLICATIONS
+Documentation Generation.
+To streamline the creation of detailed model cards, engineers can leverage model lake tasks to generate a rough draft of the required documentation.
+This application is similar to discovering and annotating metadata in data lakes (Akhtar et al., 2024; Sawadogo and Darmont, 2021; Zhao et al., 2021; Korini and Bizer, 2023; Fan et al., 2023a).
+Here’s how the process might work: upon uploading a model to the model lake, state-of-the-art techniques for tasks like attribution, versioning, benchmarking, and others can automatically analyze and map the model’s relationships to other models in the model lakes.
+The engineer can review and either accept or modify these generated mappings, especially if any information appears inaccurate, creating an initial version of the model details section.
+Additionally, the engineer can assess the model’s robustness by testing its performance against relevant benchmarks, including metrics like task accuracy, environmental impact, and more.
+Based on these test results, key sections of the model card, such as intended use and performance metrics, can be auto-populated for a faster, more reliable drafting process.
+It can alert developers to model risks, as shown by Wang et al.
+(2024a), who demonstrate how model versioning helps warn downstream model users of upstream model risks.
+Auditing.
+Policymakers have recently proposed AI safety and compliance regulations (House, 2022; Tabassi, 2023, 2024) aimed at fostering more responsible and accountable AI models (Casper et al., 2024).
+The model document generation application procedure can be repurposed for auditing by creating a template questionnaire (Golpayegani et al., 2024) and using the information from the model lake to generate a draft response with proof or explanation about how a requirement is fulfilled.
+This process can incorporate privacy-related technical solutions, where insights from model lake tasks help identify vulnerabilities across related models and their successor versions.
+It can also aid in attributing sensitive data that the model may have access to, highlighting potential exposure risks.
+Data and Model Citation.
 Data citation helps stakeholders identify the source, ownership, and authorship of the data used for a particular analysis.
-It is important to use proper data citation because the structure and contents of database can evolve (Buneman et al., 2016).
-Hence, this problem is also extended to machine learning research, since a big part of model training is its dataset.
-Based on a study conducted on 13K model cards (Liang et al., 2024), researchers found that if a training section of the model card is filled out, they generally include the name of the dataset, volume and characteristics of the dataset.
-This is not sufficient since the exact version of the dataset is usually not specified and the download links to the datasets may become stale.
-Furthermore, without any standardization and automated process on how to reference these datasets, many model cards might not contain such information or could feature varied statements regarding their referencing, as exemplified in the case shared by (Buneman et al., 2016).
-In addition to citing datasets used during the model’s training and evaluation stages, we also need to consider how to manage data generated by these models.
-It is important to do so because users can use the generated information to train another model.
-One of the solutions for at least recognizing a generated data output is use watermarks (Kirchenbauer et al., 2023).
-Another scenario is when data is generated by a model and is being used by a consumer.
-It is important to know the model’s version and input prompts to the model to improve the chances of recreating this dataset.
-Hence, there is a need to advance data citation research for models to build an ecosystem where crucial information about related databases are recorded in both data and model lakes.
-Provenance.
-A general research question for model provenance can be as follows: Given a potential source model, , and a single target model, , is the source model of ?
-If so, how is transformed into ?
-These transformations could be related to dataset being updated, training algorithms being changed, and/or changes in requirements of the model indicated by the developer as described by Mei et al.
-(2022).
-Inspired from data provenance, we can also explore where-, how- and why-provenance with respect to models.
-To get a comprehensive understanding of model provenance, it is important to know what data was used to train it.
-Using data citation helps us collect this information effectively.
-There are, however, alternative methods (Liu et al., 2023) that rely on intermediate model training checkpoints and not training data to at least verify model ownership.
-In terms of exploring why-provenance, we can try to extract datasets used or changed with the source and target models.
-For instance, there are a range of facts that can be edited using model editing techniques.
-A 1-fact edited model would mean that there was one fact used whose association is different from its source model.
-For instance, if ROME algorithm (Meng et al., 2022) was used to change the location of the Eiffel Tower from Paris to Rome, we intend to edit one sentence representation from “Eiffel Tower is located in Paris” to “Eiffel Tower is located in Rome”.
-Hence, the resulting model would be a 1-fact edited model of its source model.
-The first comparison that can be made within this sub-area, i.e., can we determine whether one model is 1-fact edited model of another.
-A general approach to understanding such models would be to reverse engineer the model’s prediction of outputs, which can lead to circuit discovery (Wang et al., 2022; Conmy et al., 2023; Elhage et al., 2021) or even understanding how to unlearn knowledge learned in models (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2023; Belrose et al., 2023; Kumari et al., 2023).
-We can extend the idea of such reverse engineering techniques to understand the process of 1-fact editing algorithms such as ROME on such models.
-Once such an ”edit or change” detector can be created, we could modify and extend its application.
-From model editing, we can then move on to model partial-tuning where a subset of the models’ parameters are used for training.
-We can look into the changed values for those parameters and understand what changed between its source and its current form.
-After building detectors for such cases, we can extend them further to understand the difference between one model that was fine-tuned from another model.
-Within fine-tuning, there are special cases of one-shot and few-shot training.
-Understanding these cases could help us build better detectors for the completely fine-tuned models where all model parameter values are usually changed.
-By creating such detectors, we can explain the transformation between a source and a target model, which is required for an exhaustive understanding of the target model’s provenance.
-Model Version Search.
-While model provenance is a research challenge for identifying and explaining the transformation of one model, an extension to this challenge could be to identify versions given a set of models, and a set of datasets, .
-A model is considered a version of another when one can be transformed into another (such as through updates in dataset and/or training algorithm) and have the same true source model (i.e., the same first version).
-Apart from the approaches mentioned earlier, an alternative way to explore versioning is to understand how different entities have similar and/or different information present in various models.
-This could, for instance, be a post-hoc analysis through creation of knowledge graphs based on particular facts in data lakes that we know change from time to time, such as the President of the United States.
-By doing so, we could create a version history for data through models published at various time period, and by extension, identify the chronological order of the models based on their knowledge space.
-Finally, by being able to pursue any direction indicated in the roadmap above, we can create a dataset or benchmark for such research so that the research community can build upon and create better methods for detecting such model differences.
-In addition to such benchmarks, it is also crucial to create metrics to understand how well these methods perform.
-By doing so, we can also increase outside-the-box access of AI models, where obtaining verified contextual information about the model can provide better transparency and evaluation for AI audits (Casper et al., 2024).
-Lastly, the solution to these methods would need visualization mechanisms to illustrate how these models and data are related.
-For instance, maybe we can have a hierarchical/network map of how different models are connected together?
-The next set of questions would be which parts of this mechanism can be automated and which ones would require a human-in-the-loop.
-5. Granularity of Model Differences
-There are various levels of granularity of model differences to consider.
-One aspect of granularity is explaining the transformation of a model.
-A low level granularity to identify the difference would be to identify the set of facts or data that were changed or were affected.
-A medium level of granularity could look like identifying the parameters that were changed and what dataset was it most likely trained on.
-A high level of granularity would be to just identify whether a model-in-question was fine-tuned from other model.
-Another aspect of granularity is the level of information used to understand the model differences – did we use just the model weights?
-do we need their datasets?
-did we also use their documentation, i.e., their model cards?
-If we can find the difference between models by just using their model weights, then we can consider this difference to be a low-level information granularity.
-If we assume that we know at least one of the models’ datasets, then we can consider the difference to be a medium level.
-Lastly, if we depend on documentation to understand a pair of models’ differences, then we would consider it as a high-level of granularity.
-This can also scale the level of trust or confidence we can have for each model difference or similarity identification.
-For instance, identifying the difference between two models using just documentation has low level of confidence because we rely on trusting that the documentation is accurate.
-Furthermore, the accessibility of documentation can be limited because there is currently no structured way of documenting a model’s training process and inputs.
-However, if we rely on just the model’s weights, we can trust the identification more since it does not have an extra layer of information to process and understand the model.
-Caveats and Considerations.
-There are a couple concerns with respect to extracting information about a model with low level granularity.
-If we were able to extract specific training data facts using just model weights, it can also be considered a privacy attack, which would mean that the model is unsafe in terms of privacy.
-Hence, in terms of providing granularity/verification scores as shown in Figure 1, it would have a high data verification score, but a low confidence in privacy.
-Another challenge with model lake tasks is differentiating between a predictive path and a deterministic path.
-For instance, in the case of identifying the source model of a certain model (model provenance task), we might find a couple of models that meet the possibility of it being a source model.
-In such cases, a lower than normal granularity score can be assigned unless a potential source model option is backed by another source of information, such as the model card.
-6. MODEL LAKES: APPLICATION
-Similar to a data lake, a model lake can be defined as a system or repository of many heterogeneous models and their related data stored in their natural format, for example, as files of model parameters.
-As the community progresses along the proposed directions, we can develop an intuitive navigation of a model lake, as depicted in Figure 1.
-This improvement can be helpful for addressing the Model Selection Problem shared in Example 1.
-In the left-side of Figure 1, we show that a user can search for models by typing out the task and type of data they are interested in working with.
-Based on this input, there can be an algorithm to semantically parse this text and find closely related task and data tags.
-Based on these tags, the search engine can share the most relevant models.
-Upon clicking a model, the model lake interface can lead to a model card page that may look similar to the right-side of the same figure.
-In this page, the users can view the model card description as well as what the model’s previous, next, source, and related models.
-In addition, they can also view what the training data was in order to understand if it was trained on desired text.
-If the training data is not specified, another way of understanding whether the model has a semantic understanding of the type of text the user cares about (for instance, legal jargon in the case of Example 1) is by viewing a semantic knowledge graph of the model and filtering for ‘legal’ within that graph.
-Within the model card, there can also be a privacy scale of low, medium, and high to gauge whether the model tends to share private information that was used in training data.
-Lastly, it could also have a set of verification scores (verifiable documentation task) to gauge how well this model is verified.
-The verification is categorized into data, code, and document verification.
-This scale would vary based on the level of granularity that they were verified with.
-0 would indicate that they haven’t been verified and 1-5 would represent high level to low level verification where where high level implies less aspects of the model verified verses low level indicates that most, if not all, the verification was done using the model weight values.
+It is important to use proper data citation because the structure and contents of the database can evolve (Buneman et al., 2016).
+Hence, this problem is also extended to machine learning, since a large part of model training is its dataset.
+Thus, the task of citing data for data lakes remains relevant for documenting the training data used in the creation of the model.
+Similarly, model citation is essential, as users can further train the model or use its outputs for consumption or additional training.
+One proposed solution to identify generated output is the use of watermarks (Kirchenbauer et al., 2023).
+In addition, model versioning tasks provide crucial documentation, allowing researchers, engineers, and developers to refer to the exact version of the model used for training or content generation.
+If a particular model is used, the platform would refer to its versioning graph and generate a citation with the model version and timestamp of the graph.
+Upon any updates of the graph, a new citation would be generated with the updated version and timestamp.
+This would be useful for also citing any generated content from this model.
+Model Search.
+In Example 1.1, we illustrate a potential scenario where a user seeks the most suitable model for summarizing and simplifying legal documents.
+As the model search task within the model lake evolves, we aim for users to be able to write declarative queries and retrieve a set of models ranked by their suitability for the specified task.
+Query example include "Find all models trained on this corpus of US Supreme Court cases" or "Find models that out perform Model X on Benchmark Y".
+Given a search task, the model lake framework can map the task function to a suitable indexer and run that indexer to retrieve top-ranked models.
+Whether the user is a technical or non-technical consumer, researcher, or engineer, they would be able to click on a model to view its model card.
+Attribution reveals how training data or learned concepts influence outputs, while version management clarifies the model’s training process, lineage, and differences, enhancing transparency.
+Benchmarking evaluates the model’s robustness on related tasks, addressing completeness.
+Together, these allow users to make informed model choices.
 7. Conclusion
-The database community has responded to the “Big Model” revolution by proposing platforms like Agora (Traub et al., 2020), that manage data-related assets, including models, datasets, software, and compute resources in a coherent ecosystem.
-There is now a need for general methods for managing model lakes.
-It will be important that the methods generalize, irrespective of how many models we are trying to understand, how they are trained, or what base of information we wish to use to resolve our tasks.
-We call on the database community to contribute to the vision of model lakes, supporting users to better distinguish a set of models that they might want to choose from for their own tasks as well as alleviate future data and model management concerns that come up the rise of deep learning models.
-Our vision is for a fundamentally new platform that extends and integrates works on data/model provenance, data/model citation, and data/model version management.
+The database community has responded to the “Big Model" revolution by proposing platforms like Agora (Traub et al., 2020), that manage data-related assets, including models, datasets, software, and compute resources in a coherent ecosystem.
+But the model-specific support in such an ecosystem must be expanded to include general methods for managing and finding models, support that can make such systems fully functional model lakes.
+It will be important that the methods generalize, irrespective of how many models we are trying to understand, what architectures the models use, how they are trained, or how we wish to query or search the models.
+We call on the database community to contribute to the vision of model lakes, supporting users to more easily find relevant models and to better understand those models.
+Our vision is for a fundamentally new platform that extends and integrates work on data/model attribution, data/model search, and data/model version management.
 Acknowledgements.
-KP and RM was supported in part by NSF award numbers IIS-2107248, IIS-1956096, IIS-2325632, and KP and DB by a grant from Open Philanthropy.References
+KP and RM were supported in part by NSF award numbers IIS-2107248, IIS-1956096, IIS-2325632, and KP and DB by a grant from Open Philanthropy.References
 - (1)
+- Achille et al.
+(2019) Alessandro Achille, Michael Lam, Rahul Tewari, Avinash Ravichandran, Subhransu Maji, Charless Fowlkes, Stefano Soatto, and Pietro Perona.
+2019.
+Task2Vec: Task Embedding for Meta-Learning.
+In 2019 IEEE/CVF International Conference on Computer Vision (ICCV).
+6429–6438.
+https://doi.org/10.1109/ICCV.2019.00653
 - AI ([n.d.]a) ELSA AI. [n.d.]a. ELSA. https://elsaspeak.com/en/.
 - AI ([n.d.]b) Fitness AI. [n.d.]b. Fitness AI. https://www.fitnessai.com/.
-- Alrashed et al.
-(2021) Tarfah Alrashed, Dimitris Paparas, Omar Benjelloun, Ying Sheng, and Natasha F.
-Noy.
-2021.
-Dataset or Not?
-A Study on the Veracity of Semantic Markup for Dataset Pages.
-In The Semantic Web - ISWC 2021 - 20th International Semantic Web Conference, ISWC 2021, Virtual Event, October 24-28, 2021, Proceedings (Lecture Notes in Computer Science), Andreas Hotho, Eva Blomqvist, Stefan Dietze, Achille Fokoue, Ying Ding, Payam M.
-Barnaghi, Armin Haller, Mauro Dragoni, and Harith Alani (Eds.), Vol.
-12922.
-Springer, 338–356.
-https://doi.org/10.1007/978-3-030-88361-4_20
+- Akhtar et al.
+(2024) Mubashara Akhtar, Omar Benjelloun, Costanza Conforti, Pieter Gijsbers, Joan Giner-Miguelez, Nitisha Jain, Michael Kuchnik, Quentin Lhoest, Pierre Marcenac, Manil Maskey, Peter Mattson, Luis Oala, Pierre Ruyssen, Rajat Shinde, Elena Simperl, Goeffry Thomas, Slava Tykhonov, Joaquin Vanschoren, Jos van der Velde, Steffen Vogler, and Carole-Jean Wu.
+2024.
+Croissant: A Metadata Format for ML-Ready Datasets.
+In Proceedings of the Eighth Workshop on Data Management for End-to-End Machine Learning (DEEM ’24).
+Association for Computing Machinery, New York, NY, USA, 1–6.
+https://doi.org/10.1145/3650203.3663326
+- Andy Zou (2023) Sarah Chen James Campbell Phillip Guo Richard Ren Alexander Pan Xuwang Yin Mantas Mazeika Ann-Kathrin Dombrowski Shashwat Goel Nathaniel Li Michael J.
+Byun Zifan Wang Alex Mallen Steven Basart Sanmi Koyejo Dawn Song Matt Fredrikson Zico Kolter Dan Hendrycks Andy Zou, Long Phan.
+2023.
+Representation Engineering: A Top-Down Approach to AI Transparency.
+arXiv:cs.CL/2310.01405
+- Anil et al.
+(2023) Rohan Anil, Sebastian Borgeaud, Yonghui Wu, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M.
+Dai, Anja Hauth, Katie Millican, David Silver, Slav Petrov, Melvin Johnson, Ioannis Antonoglou, Julian Schrittwieser, Amelia Glaese, Jilin Chen, Emily Pitler, Timothy P.
+Lillicrap, Angeliki Lazaridou, Orhan Firat, James Molloy, Michael Isard, Paul Ronald Barham, Tom Hennigan, Benjamin Lee, Fabio Viola, Malcolm Reynolds, Yuanzhong Xu, Ryan Doherty, Eli Collins, Clemens Meyer, Eliza Rutherford, Erica Moreira, Kareem Ayoub, Megha Goel, George Tucker, Enrique Piqueras, Maxim Krikun, Iain Barr, Nikolay Savinov, Ivo Danihelka, Becca Roelofs, Anaïs White, Anders Andreassen, Tamara von Glehn, Lakshman Yagati, Mehran Kazemi, Lucas Gonzalez, Misha Khalman, Jakub Sygnowski, and et al.
+2023.
+Gemini: A Family of Highly Capable Multimodal Models.
+CoRR abs/2312.11805 (2023).
+https://doi.org/10.48550/ARXIV.2312.11805 arXiv:2312.11805
+- Anthropic (2024) Anthropic.
+2024.
+The Claude 3 Model Family: Opus, Sonnet, Haiku.
+https://www-cdn.anthropic.com/de8ba9b01c9ab7cbabf5c33b80b7bbc618857627/Model_Card_Claude_3.pdf.
+Accessed: 2024-09-30.
 - Arnold et al.
@@
 LEACE: Perfect linear concept erasure in closed form.
-arXiv:2306.03819 [cs.LG]
-- Benjelloun et al.
-(2020) Omar Benjelloun, Shiyu Chen, and Natasha F.
-Noy.
-2020.
-Google Dataset Search by the Numbers.
-In The Semantic Web - ISWC 2020 - 19th International Semantic Web Conference, Athens, Greece, November 2-6, 2020, Proceedings, Part II (Lecture Notes in Computer Science), Jeff Z.
-Pan, Valentina A.
-M.
-Tamma, Claudia d’Amato, Krzysztof Janowicz, Bo Fu, Axel Polleres, Oshani Seneviratne, and Lalana Kagal (Eds.), Vol.
-12507.
-Springer, 667–682.
-https://doi.org/10.1007/978-3-030-62466-8_41
-- Bernstein et al.
-(2000) Philip A.
-Bernstein, Alon Y.
-Halevy, and Rachel Pottinger.
-2000.
-A Vision of Management of Complex Models.
-SIGMOD Rec.
-29, 4 (2000), 55–63.
-https://doi.org/10.1145/369275.369289
-- Bhardwaj et al.
-(2015) Anant P.
-Bhardwaj, Amol Deshpande, Aaron J.
-Elmore, David R.
-Karger, Sam Madden, Aditya G.
-Parameswaran, Harihar Subramanyam, Eugene Wu, and Rebecca Zhang.
-2015.
-Collaborative Data Analytics with DataHub.
-Proc.
-VLDB Endow.
-8, 12 (2015), 1916–1919.
-https://doi.org/10.14778/2824032.2824100
-- Borisov et al.
-(2022) Vadim Borisov, Tobias Leemann, Kathrin Seßler, Johannes Haug, Martin Pawelczyk, and Gjergji Kasneci.
-2022.
-Deep Neural Networks and Tabular Data: A Survey.
-IEEE Transactions on Neural Networks and Learning Systems (2022), 1–21.
-https://doi.org/10.1109/TNNLS.2022.3229161
-- Brown et al.
-(2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al.
-2020.
-Language models are few-shot learners.
-Advances in neural information processing systems 33 (2020), 1877–1901.
+arXiv:cs.LG/2306.03819
+- Bernstein and Melnik (2007) Philip A.
+Bernstein and Sergey Melnik.
+2007.
+Model management 2.0: manipulating richer mappings.
+In Proceedings of the 2007 ACM SIGMOD International Conference on Management of Data (SIGMOD ’07).
+Association for Computing Machinery, New York, NY, USA, 1–12.
+https://doi.org/10.1145/1247480.1247482
+- Bogatu et al. (2020) Alex Bogatu, Alvaro A. A. Fernandes, Norman W. Paton, and Nikolaos Konstantinou. 2020. Dataset Discovery in Data Lakes. In ICDE. 709–720.
 - Buneman et al.
@@
 - Buneman et al. (2016) Peter Buneman, Susan B. Davidson, and James Frew. 2016. Why data citation is a computational problem. Commun. ACM 59, 9 (2016), 50–57. https://doi.org/10.1145/2893181
+- Buneman et al.
+(2001) Peter Buneman, Sanjeev Khanna, and Tan Wang-Chiew.
+2001.
+Why and where: A characterization of data provenance.
+In Database Theory—ICDT 2001: 8th International Conference London, UK, January 4–6, 2001 Proceedings 8.
+Springer, 316–330.
 - Buneman and Tan (2018) Peter Buneman and Wang-Chiew Tan. 2018. Data Provenance: What next? SIGMOD Rec. 47, 3 (2018), 5–16. https://doi.org/10.1145/3316416.3316418
+- Carlini et al.
+(2021) Nicholas Carlini, Florian Tramer, Eric Wallace, Matthew Jagielski, Ariel Herbert-Voss, Katherine Lee, Adam Roberts, Tom Brown, Dawn Song, Ulfar Erlingsson, et al.
+2021.
+Extracting training data from large language models.
+In 30th USENIX Security Symposium (USENIX Security 21).
+2633–2650.
 - Casper et al.
@@
 Black-Box Access is Insufficient for Rigorous AI Audits.
-arXiv:2401.14446 [cs.CY]
-- Chapman et al.
-(2020) Adriane Chapman, Elena Simperl, Laura Koesten, George Konstantinidis, Luis-Daniel Ibáñez, Emilia Kacprzak, and Paul Groth.
-2020.
-Dataset search: a survey.
-VLDB J.
-29, 1 (2020), 251–272.
-https://doi.org/10.1007/S00778-019-00564-X
+arXiv:cs.CY/2401.14446
 - Chen et al. (2023) Hugh Chen, Ian C Covert, Scott M Lundberg, and Su-In Lee. 2023. Algorithms to estimate Shapley value feature attributions. Nature Machine Intelligence (2023), 1–12.
@@
 PMLR, 354–372.
-- Chung et al.
-(2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, Albert Webson, Shixiang Shane Gu, Zhuyun Dai, Mirac Suzgun, Xinyun Chen, Aakanksha Chowdhery, Alex Castro-Ros, Marie Pellat, Kevin Robinson, Dasha Valter, Sharan Narang, Gaurav Mishra, Adams Yu, Vincent Zhao, Yanping Huang, Andrew Dai, Hongkun Yu, Slav Petrov, Ed H.
-Chi, Jeff Dean, Jacob Devlin, Adam Roberts, Denny Zhou, Quoc V.
-Le, and Jason Wei.
-2022.
-Scaling Instruction-Finetuned Language Models.
-arXiv:2210.11416 [cs.LG]
 - Clark et al.
@@
 Towards Automated Circuit Discovery for Mechanistic Interpretability.
-arXiv:2304.14997 [cs.LG]
-- Cun et al.
-(1990) Y.
-Le Cun, B.
-Boser, J.
+arXiv:cs.LG/2304.14997
+- DeepSeek-AI et al.
+(2025) DeepSeek-AI, Daya Guo, Dejian Yang, Haowei Zhang, Junxiao Song, Ruoyu Zhang, Runxin Xu, Qihao Zhu, Shirong Ma, Peiyi Wang, Xiao Bi, Xiaokang Zhang, Xingkai Yu, Yu Wu, Z.
+F.
+Wu, Zhibin Gou, Zhihong Shao, Zhuoshu Li, Ziyi Gao, Aixin Liu, Bing Xue, Bingxuan Wang, Bochao Wu, Bei Feng, Chengda Lu, Chenggang Zhao, Chengqi Deng, Chenyu Zhang, Chong Ruan, Damai Dai, Deli Chen, Dongjie Ji, Erhang Li, Fangyun Lin, Fucong Dai, Fuli Luo, Guangbo Hao, Guanting Chen, Guowei Li, H.
+Zhang, Han Bao, Hanwei Xu, Haocheng Wang, Honghui Ding, Huajian Xin, Huazuo Gao, Hui Qu, Hui Li, Jianzhong Guo, Jiashi Li, Jiawei Wang, Jingchang Chen, Jingyang Yuan, Junjie Qiu, Junlong Li, J.
+L.
+Cai, Jiaqi Ni, Jian Liang, Jin Chen, Kai Dong, Kai Hu, Kaige Gao, Kang Guan, Kexin Huang, Kuai Yu, Lean Wang, Lecong Zhang, Liang Zhao, Litong Wang, Liyue Zhang, Lei Xu, Leyi Xia, Mingchuan Zhang, Minghua Zhang, Minghui Tang, Meng Li, Miaojun Wang, Mingming Li, Ning Tian, Panpan Huang, Peng Zhang, Qiancheng Wang, Qinyu Chen, Qiushi Du, Ruiqi Ge, Ruisong Zhang, Ruizhe Pan, Runji Wang, R.
+J.
+Chen, R.
+L.
+Jin, Ruyi Chen, Shanghao Lu, Shangyan Zhou, Shanhuang Chen, Shengfeng Ye, Shiyu Wang, Shuiping Yu, Shunfeng Zhou, Shuting Pan, S.
 S.
-Denker, R.
-E.
-Howard, W.
-Habbard, L.
-D.
-Jackel, and D.
-Henderson.
-1990.
-Handwritten digit recognition with a back-propagation network.
-Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 396–404.
-- Dai et al.
-(2022) Damai Dai, Li Dong, Yaru Hao, Zhifang Sui, Baobao Chang, and Furu Wei.
-2022.
-Knowledge Neurons in Pretrained Transformers.
-In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers).
-Association for Computational Linguistics, Dublin, Ireland, 8493–8502.
-https://doi.org/10.18653/v1/2022.acl-long.581
-- De Cao et al.
-(2021) Nicola De Cao, Wilker Aziz, and Ivan Titov.
-2021.
-Editing Factual Knowledge in Language Models.
-In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing.
-Association for Computational Linguistics, Online and Punta Cana, Dominican Republic, 6491–6506.
-https://doi.org/10.18653/v1/2021.emnlp-main.522
+Li, Shuang Zhou, Shaoqing Wu, Shengfeng Ye, Tao Yun, Tian Pei, Tianyu Sun, T.
+Wang, Wangding Zeng, Wanjia Zhao, Wen Liu, Wenfeng Liang, Wenjun Gao, Wenqin Yu, Wentao Zhang, W.
+L.
+Xiao, Wei An, Xiaodong Liu, Xiaohan Wang, Xiaokang Chen, Xiaotao Nie, Xin Cheng, Xin Liu, Xin Xie, Xingchao Liu, Xinyu Yang, Xinyuan Li, Xuecheng Su, Xuheng Lin, X.
+Q.
+Li, Xiangyue Jin, Xiaojin Shen, Xiaosha Chen, Xiaowen Sun, Xiaoxiang Wang, Xinnan Song, Xinyi Zhou, Xianzu Wang, Xinxia Shan, Y.
+K.
+Li, Y.
+Q.
+Wang, Y.
+X.
+Wei, Yang Zhang, Yanhong Xu, Yao Li, Yao Zhao, Yaofeng Sun, Yaohui Wang, Yi Yu, Yichao Zhang, Yifan Shi, Yiliang Xiong, Ying He, Yishi Piao, Yisong Wang, Yixuan Tan, Yiyang Ma, Yiyuan Liu, Yongqiang Guo, Yuan Ou, Yuduan Wang, Yue Gong, Yuheng Zou, Yujia He, Yunfan Xiong, Yuxiang Luo, Yuxiang You, Yuxuan Liu, Yuyang Zhou, Y.
+X.
+Zhu, Yanhong Xu, Yanping Huang, Yaohui Li, Yi Zheng, Yuchen Zhu, Yunxian Ma, Ying Tang, Yukun Zha, Yuting Yan, Z.
+Z.
+Ren, Zehui Ren, Zhangli Sha, Zhe Fu, Zhean Xu, Zhenda Xie, Zhengyan Zhang, Zhewen Hao, Zhicheng Ma, Zhigang Yan, Zhiyu Wu, Zihui Gu, Zijia Zhu, Zijun Liu, Zilin Li, Ziwei Xie, Ziyang Song, Zizheng Pan, Zhen Huang, Zhipeng Xu, Zhongyu Zhang, and Zhen Zhang.
+2025.
+DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning.
+arXiv:cs.CL/2501.12948 https://arxiv.org/abs/2501.12948
+- Deng et al.
+(2023) Xiang Deng, Yu Gu, Boyuan Zheng, Shijie Chen, Samuel Stevens, Boshi Wang, Huan Sun, and Yu Su.
+2023.
+Mind2Web: Towards a Generalist Agent for the Web.
+In Thirty-seventh Conference on Neural Information Processing Systems.
+https://openreview.net/forum?id=kiYqbO3wqw
 - Devlin et al.
@@
 LIDA: A Tool for Automatic Generation of Grammar-Agnostic Visualizations and Infographics using Large Language Models.
-In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics: System Demonstrations, ACL 2023, Toronto, Canada, July 10-12, 2023, Danushka Bollegala, Ruihong Huang, and Alan Ritter (Eds.).
-Association for Computational Linguistics, 113–126.
+In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 3: System Demonstrations), Danushka Bollegala, Ruihong Huang, and Alan Ritter (Eds.).
+Association for Computational Linguistics, Toronto, Canada, 113–126.
 https://doi.org/10.18653/v1/2023.acl-demo.11
-- Doan et al. (2012) AnHai Doan, Alon Y. Halevy, and Zachary G. Ives. 2012. Principles of Data Integration. Morgan Kaufmann. http://research.cs.wisc.edu/dibook/
+- Dong et al.
+(2022) Qingxiu Dong, Lei Li, Damai Dai, Ce Zheng, Zhiyong Wu, Baobao Chang, Xu Sun, Jingjing Xu, and Zhifang Sui.
+2022.
+A Survey for In-context Learning.
+arXiv preprint arXiv:2301.00234 (2022).
 - Dong et al.
@@
 - Doshi-Velez and Kim (2017) Finale Doshi-Velez and Been Kim. 2017. Towards a rigorous science of interpretable machine learning. arXiv preprint arXiv:1702.08608 (2017).
-- Eldan and Russinovich (2023) Ronen Eldan and Mark Russinovich. 2023. Who’s Harry Potter? Approximate Unlearning in LLMs. arXiv:2310.02238 [cs.CL]
+- Dou et al.
+(2024) Yao Dou, Isadora Krsek, Tarek Naous, Anubha Kabra, Sauvik Das, Alan Ritter, and Wei Xu.
+2024.
+Reducing Privacy Risks in Online Self-Disclosures with Language Models.
+In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.).
+Association for Computational Linguistics, Bangkok, Thailand, 13732–13754.
+https://doi.org/10.18653/v1/2024.acl-long.741
+- Dwork et al.
+(2006) Cynthia Dwork, Frank McSherry, Kobbi Nissim, and Adam Smith.
+2006.
+Calibrating noise to sensitivity in private data analysis.
+In Theory of Cryptography: Third Theory of Cryptography Conference, TCC 2006, New York, NY, USA, March 4-7, 2006.
+Proceedings 3.
+Springer, 265–284.
+- Eilertsen et al.
+(2020) Gabriel Eilertsen, Daniel Jönsson, Timo Ropinski, Jonas Unger, and Anders Ynnerman.
+2020.
+Classifying the classifier: dissecting the weight space of neural networks.
+In ECAI 2020.
+IOS Press, 1119–1126.
+- Eldan and Russinovich (2023) Ronen Eldan and Mark Russinovich. 2023. Who’s Harry Potter? Approximate Unlearning in LLMs. arXiv:cs.CL/2310.02238
 - Elhage et al.
@@
 Table Discovery in Data Lakes: State-of-the-art and Future Directions.
-In Companion of the 2023 International Conference on Management of Data (Seattle, WA, USA) (SIGMOD ’23).
+In Companion of the 2023 International Conference on Management of Data (SIGMOD ’23).
 Association for Computing Machinery, New York, NY, USA, 69–75.
@@
 https://doi.org/10.1109/DSAA.2018.00018
+- Golpayegani et al.
+(2024) Delaram Golpayegani, Isabelle Hupont, Cecilia Panigutti, Harshvardhan J Pandit, Sven Schade, Declan O’Sullivan, and Dave Lewis.
+2024.
+AI cards: towards an applied framework for machine-readable AI and risk documentation inspired by the EU AI Act.
+In Annual Privacy Forum.
+Springer, 48–72.
 - Google ([n.d.]a) Google. [n.d.]a. BARD. https://bard.google.com/chat.
 - Google ([n.d.]b) Google. [n.d.]b. Socratic. https://socratic.org/.
-- Ha et al. (2017) David Ha, Andrew M. Dai, and Quoc V. Le. 2017. HyperNetworks. In International Conference on Learning Representations. https://openreview.net/forum?id=rkpACe1lx
-- Halevy et al.
-(2023) Alon Y.
-Halevy, Yejin Choi, Avrilia Floratou, Michael J.
-Franklin, Natasha F.
-Noy, and Haixun Wang.
+- Goyal et al.
+(2017) Yash Goyal, Tejas Khot, Douglas Summers-Stay, Dhruv Batra, and Devi Parikh.
+2017.
+Making the v in VQA Matter: Elevating the Role of Image Understanding in Visual Question Answering.
+In Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition (CVPR).
+- Grosse et al.
+(2023) Roger Grosse, Juhan Bae, Cem Anil, Nelson Elhage, Alex Tamkin, Amirhossein Tajdini, Benoit Steiner, Dustin Li, Esin Durmus, Ethan Perez, et al.
 2023.
-Will LLMs reshape, supercharge, or kill data science?
-Proc.
-VLDB Endow.
-16, 12 (2023), 4114–4115.
-https://doi.org/10.14778/3611540.3611634
+Studying large language model generalization with influence functions.
+arXiv preprint arXiv:2308.03296 (2023).
+- Hendrycks et al.
+(2021a) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt.
+2021a.
+Measuring Massive Multitask Language Understanding.
+In 9th International Conference on Learning Representations, ICLR 2021, Virtual Event, Austria, May 3-7, 2021.
+OpenReview.net.
+https://openreview.net/forum?id=d7KBjmI3GmQ
+- Hendrycks et al.
+(2021b) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt.
+2021b.
+Measuring Mathematical Problem Solving With the MATH Dataset.
+In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 2).
+https://openreview.net/forum?id=7Bywt2mQsCe
+- Heusel et al.
+(2017) Martin Heusel, Hubert Ramsauer, Thomas Unterthiner, Bernhard Nessler, and Sepp Hochreiter.
+2017.
+Gans trained by a two time-scale update rule converge to a local nash equilibrium.
+Advances in neural information processing systems 30 (2017).
+- Horwitz et al. (2024) Eliahu Horwitz, Asaf Shul, and Yedid Hoshen. 2024. On the Origin of Llamas: Model Tree Heritage Recovery. arXiv preprint arXiv:2405.18432 (2024).
+- House (2022) The White House. 2022. Blueprint for an AI Bill of Rights: Making Automated Systems Work for the American People. The White House.
 - Hu et al.
@@
 https://openreview.net/forum?id=nZeVKeeFYf9
+- Hu et al.
+(2023) Xuming Hu, Shen Wang, Xiao Qin, Chuan Lei, Zhengyuan Shen, Christos Faloutsos, Asterios Katsifodimos, George Karypis, Lijie Wen, and Philip S.
+Yu.
+2023.
+Automatic Table Union Search with Tabular Representation Learning.
+In Findings of the Association for Computational Linguistics: ACL 2023, Toronto, Canada, July 9-14, 2023.
+Association for Computational Linguistics, 3786–3800.
+https://aclanthology.org/2023.findings-acl.233
+- Huang et al.
+(2024) Xinting Huang, Madhur Panwar, Navin Goyal, and Michael Hahn.
+2024.
+InversionView: A General-Purpose Method for Reading Information from Neural Activations.
+arXiv:cs.LG/2405.17653 https://arxiv.org/abs/2405.17653
 - Inc. (2023) Civit AI Inc. 2023. https://civitai.com/
@@
 https://doi.org/10.18653/v1/2023.acl-long.805
+- Jiang et al.
+(2023) Albert Q.
+Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, Diego de Las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, Lélio Renard Lavaud, Marie-Anne Lachaux, Pierre Stock, Teven Le Scao, Thibaut Lavril, Thomas Wang, Timothée Lacroix, and William El Sayed.
+2023.
+Mistral 7B.
+CoRR abs/2310.06825 (2023).
+https://doi.org/10.48550/ARXIV.2310.06825 arXiv:2310.06825
 - Johnson et al.
@@
 - Khatiwada et al. (2022) Aamod Khatiwada, Roee Shraga, Wolfgang Gatterbauer, and Renée J Miller. 2022. Integrating Data Lake Tables. Proceedings of the VLDB Endowment 16, 4 (2022), 932–945.
-- Kingma and Ba (2015) Diederik P.
-Kingma and Jimmy Ba.
-2015.
-Adam: A Method for Stochastic Optimization.
-In 3rd International Conference on Learning Representations, ICLR 2015, San Diego, CA, USA, May 7-9, 2015, Conference Track Proceedings, Yoshua Bengio and Yann LeCun (Eds.).
-http://arxiv.org/abs/1412.6980
 - Kirchenbauer et al.
@@
 https://proceedings.mlr.press/v202/kirchenbauer23a.html
+- Koh and Liang (2017) Pang Wei Koh and Percy Liang. 2017. Understanding black-box predictions via influence functions. In International conference on machine learning. PMLR, 1885–1894.
 - Korini and Bizer (2023) Keti Korini and Christian Bizer.
@@
 In International Conference on Computer Vision (ICCV).
-- Labs ([n.d.]a) Neptune Labs. [n.d.]a. neptune.ai — The MLOps stack component for experiment tracking. https://neptune.ai/.
+- Labs ([n.d.]a) Neptune Labs. [n.d.]a. neptune.ai | The MLOps stack component for experiment tracking. https://neptune.ai/.
 - Labs ([n.d.]b) Prisma Labs. [n.d.]b. Lensa. https://prisma-ai.com/lensa.
+- Lacoste et al. (2019) Alexandre Lacoste, Alexandra Luccioni, Victor Schmidt, and Thomas Dandres. 2019. Quantifying the Carbon Emissions of Machine Learning. arXiv preprint arXiv:1910.09700 (2019).
+- Lenc and Vedaldi (2015) Karel Lenc and Andrea Vedaldi.
+2015.
+Understanding image representations by measuring their equivariance and equivalence.
+In Proceedings of the IEEE conference on computer vision and pattern recognition.
+991–999.
 - Lewis et al.
@@
 - Li et al.
-(2020) Yuliang Li, Jinfeng Li, Yoshihiko Suhara, AnHai Doan, and Wang-Chiew Tan.
-2020.
-Deep Entity Matching with Pre-Trained Language Models.
-14, 1 (sep 2020), 50–60.
-https://doi.org/10.14778/3421424.3421431
+(2025) Tianlong Li, Zhenghua Wang, Wenhao Liu, Muling Wu, Shihan Dou, Changze Lv, Xiaohua Wang, Xiaoqing Zheng, and Xuanjing Huang.
+2025.
+Revisiting Jailbreaking for Large Language Models: A Representation Engineering Perspective.
+In Proceedings of the 31st International Conference on Computational Linguistics, Owen Rambow, Leo Wanner, Marianna Apidianaki, Hend Al-Khalifa, Barbara Di Eugenio, and Steven Schockaert (Eds.).
+Association for Computational Linguistics, Abu Dhabi, UAE, 3158–3178.
+https://aclanthology.org/2025.coling-main.212/
 - Li et al.
@@
 Systematic Analysis of 32K AI Model Cards.
-arXiv:2402.05160 [cs.SE]
+arXiv:cs.SE/2402.05160
+- Lin et al.
+(2014) Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva Ramanan, Piotr Dollár, and C Lawrence Zitnick.
+2014.
+Microsoft coco: Common objects in context.
+In Computer Vision–ECCV 2014: 13th European Conference, Zurich, Switzerland, September 6-12, 2014, Proceedings, Part V 13.
+Springer, 740–755.
 - Liu et al.
-(2023) Yunpeng Liu, Kexin Li, Zhuotao Liu, Bihan Wen, Ke Xu, Weiqiang Wang, Wenbiao Zhao, and Qi Li.
+(2023) Pengfei Liu, Weizhe Yuan, Jinlan Fu, Zhengbao Jiang, Hiroaki Hayashi, and Graham Neubig.
 2023.
-Provenance of Training without Training Data: Towards Privacy-Preserving DNN Model Ownership Verification.
-In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23).
-Association for Computing Machinery, New York, NY, USA, 1980–1990.
-https://doi.org/10.1145/3543507.3583198
+Pre-train, Prompt, and Predict: A Systematic Survey of Prompting Methods in Natural Language Processing.
+ACM Comput.
+Surv.
+55, 9, Article 195 (Jan.
+2023), 35 pages.
+https://doi.org/10.1145/3560815
+- Longpre et al.
+(2024) Shayne Longpre, Stella Biderman, Alon Albalak, Hailey Schoelkopf, Daniel McDuff, Sayash Kapoor, Kevin Klyman, Kyle Lo, Gabriel Ilharco, Nay San, Maribeth Rauh, Aviya Skowron, Bertie Vidgen, Laura Weidinger, Arvind Narayanan, Victor Sanh, David Ifeoluwa Adelani, Percy Liang, Rishi Bommasani, Peter Henderson, Sasha Luccioni, Yacine Jernite, and Luca Soldaini.
+2024.
+The Responsible Foundation Model Development Cheatsheet: A Review of Tools & Resources.
+CoRR abs/2406.16746 (2024).
+https://doi.org/10.48550/ARXIV.2406.16746 arXiv:2406.16746
+- Longpre et al.
+(2023) Shayne Longpre, Robert Mahari, Anthony Chen, Naana Obeng-Marnu, Damien Sileo, William Brannon, Niklas Muennighoff, Nathan Khazam, Jad Kabbara, Kartik Perisetla, Xinyi Wu, Enrico Shippole, Kurt D.
+Bollacker, Tongshuang Wu, Luis Villa, Sandy Pentland, Deb Roy, and Sara Hooker.
+2023.
+The Data Provenance Initiative: A Large Scale Audit of Dataset Licensing & Attribution in AI.
+CoRR abs/2310.16787 (2023).
+https://doi.org/10.48550/ARXIV.2310.16787 arXiv:2310.16787
 - Lu et al.
@@
 https://doi.org/10.48550/ARXIV.2401.12874 arXiv:2401.12874
+- Malkov and Yashunin (2020) Yury A.
+Malkov and Dmitry A.
+Yashunin.
+2020.
+Efficient and Robust Approximate Nearest Neighbor Search Using Hierarchical Navigable Small World Graphs.
+IEEE Trans.
+Pattern Anal.
+Mach.
+Intell.
+42, 4 (2020), 824–836.
+- Mattson et al.
+(2019) Peter Mattson, Christine Cheng, Cody Coleman, Greg Diamos, Paulius Micikevicius, David A.
+Patterson, Hanlin Tang, Gu-Yeon Wei, Peter Bailis, Victor Bittorf, David Brooks, Dehao Chen, Debojyoti Dutta, Udit Gupta, Kim M.
+Hazelwood, Andrew Hock, Xinyuan Huang, Bill Jia, Daniel Kang, David Kanter, Naveen Kumar, Jeffery Liao, Guokai Ma, Deepak Narayanan, Tayo Oguntebi, Gennady Pekhimenko, Lillian Pentecost, Vijay Janapa Reddi, Taylor Robie, Tom St.
+John, Carole-Jean Wu, Lingjie Xu, Cliff Young, and Matei Zaharia.
+2019.
+MLPerf Training Benchmark.
+CoRR abs/1910.01500 (2019).
+arXiv:1910.01500 http://arxiv.org/abs/1910.01500
 - Mei et al.
@@
 Model Provenance Management in MLOps Pipeline.
-In Proceedings of the 2022 8th International Conference on Computing and Data Engineering (Bangkok, Thailand) (ICCDE ’22).
+In Proceedings of the 2022 8th International Conference on Computing and Data Engineering (ICCDE ’22).
 Association for Computing Machinery, New York, NY, USA, 45–50.
@@
 - Microsoft ([n.d.]) Microsoft. [n.d.]. Seeing AI. https://www.microsoft.com/en-us/ai/seeing-ai.
+- Mineault et al.
+(2024) Patrick Mineault, Niccolò Zanichelli, Joanne Zichen Peng, Anton Arkhipov, Eli Bingham, Julian Jara-Ettinger, Emily Mackevicius, Adam Marblestone, Marcelo Mattar, Andrew Payne, Sophia Sanborn, Karen Schroeder, Zenna Tavares, and Andreas Tolias.
+2024.
+NeuroAI for AI Safety.
+arXiv:cs.AI/2411.18526 https://arxiv.org/abs/2411.18526
 - Mitchell et al.
@@
 Model Cards for Model Reporting.
-In Proceedings of the Conference on Fairness, Accountability, and Transparency (Atlanta, GA, USA) (FAT* ’19).
+In Proceedings of the Conference on Fairness, Accountability, and Transparency (FAT* ’19).
 Association for Computing Machinery, New York, NY, USA, 220–229.
@@
 Prov-dm: The prov data model.
-- Mu et al. (2023) Xin Mu, Yu Wang, Yehong Zhang, Jiaqi Zhang, Hui Wang, Yang Xiang, and Yue Yu. 2023. Model Provenance via Model DNA. arXiv:2308.02121 [cs.LG]
+- Morris et al.
+(2023) John Morris, Volodymyr Kuleshov, Vitaly Shmatikov, and Alexander Rush.
+2023.
+Text Embeddings Reveal (Almost) As Much As Text.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.).
+Association for Computational Linguistics, Singapore, 12448–12460.
+https://doi.org/10.18653/v1/2023.emnlp-main.765
+- Morris et al.
+(2024) John Xavier Morris, Wenting Zhao, Justin T Chiu, Vitaly Shmatikov, and Alexander M Rush.
+2024.
+Language Model Inversion.
+In The Twelfth International Conference on Learning Representations.
+https://openreview.net/forum?id=t9dWHpGkPj
+- Mu et al. (2023) Xin Mu, Yu Wang, Yehong Zhang, Jiaqi Zhang, Hui Wang, Yang Xiang, and Yue Yu. 2023. Model Provenance via Model DNA. arXiv:cs.LG/2308.02121
+- Mueller et al.
+(2024) Aaron Mueller, Jannik Brinkmann, Millicent L.
+Li, Samuel Marks, Koyena Pal, Nikhil Prakash, Can Rager, Aruna Sankaranarayanan, Arnab Sen Sharma, Jiuding Sun, Eric Todd, David Bau, and Yonatan Belinkov.
+2024.
+The Quest for the Right Mediator: A History, Survey, and Theoretical Grounding of Causal Interpretability.
+CoRR abs/2408.01416 (2024).
+https://doi.org/10.48550/ARXIV.2408.01416 arXiv:2408.01416
 - Murdoch et al.
@@
 - Nargesian et al. (2018) Fatemeh Nargesian, Erkang Zhu, Ken Q Pu, and Renée J Miller. 2018. Table union search on open data. PVLDB 11, 7 (2018), 813–825.
-- OpenAI ([n.d.]a) OpenAI. [n.d.]a. Chat GPT. https://chat.openai.com/chat.
-- OpenAI ([n.d.]b) OpenAI. [n.d.]b. Introducing GPTs. https://openai.com/blog/introducing-gpts.
+- Navon et al.
+(2023) Aviv Navon, Aviv Shamsian, Idan Achituve, Ethan Fetaya, Gal Chechik, and Haggai Maron.
+2023.
+Equivariant Architectures for Learning in Deep Weight Spaces.
+In Proceedings of the 40th International Conference on Machine Learning (Proceedings of Machine Learning Research), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.), Vol.
+202.
+PMLR, 25790–25816.
+https://proceedings.mlr.press/v202/navon23a.html
+- OpenAI ([n.d.]) OpenAI. [n.d.]. Chat GPT. https://chat.openai.com/chat.
+- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. CoRR abs/2303.08774 (2023). https://doi.org/10.48550/ARXIV.2303.08774 arXiv:2303.08774
 - Orgad et al.
@@
 https://doi.org/10.1109/ICCV51070.2023.00649
+- Ortega et al.
+(2018) Pedro A Ortega, Vishal Maini, and DeepMind Safety Team.
+2018.
+Building safe artificial intelligence: specification, robustness, and assurance.
+DeepMind Safety Research Blog (2018).
+- Ouyang et al.
+(2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L.
+Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe.
+2022.
+Training language models to follow instructions with human feedback.
+arXiv:cs.CL/2203.02155 https://arxiv.org/abs/2203.02155
 - Petsiuk et al. (2018) Vitali Petsiuk, Abir Das, and Kate Saenko. 2018. Rise: Randomized input sampling for explanation of black-box models. arXiv preprint arXiv:1806.07421 (2018).
+- Prabhu et al.
+(2024) Ameya Prabhu, Vishaal Udandarao, Philip Torr, Matthias Bethge, Adel Bibi, and Samuel Albanie.
+2024.
+Lifelong Benchmarks: Efficient Model Evaluation in an Era of Rapid Progress.
+arXiv:cs.LG/2402.19472 https://arxiv.org/abs/2402.19472
 - Raffel et al.
@@
 The Journal of Machine Learning Research 21, 1 (2020), 5485–5551.
-- Ramesh et al. (2022) Aditya Ramesh, Prafulla Dhariwal, Alex Nichol, Casey Chu, and Mark Chen. 2022. Hierarchical Text-Conditional Image Generation with CLIP Latents. arXiv:2204.06125 [cs.CV]
 - Räuker et al.
@@
 https://doi.org/10.48550/ARXIV.2207.13243 arXiv:2207.13243
+- Reddi et al.
+(2019) Vijay Janapa Reddi, Christine Cheng, David Kanter, Peter Mattson, Guenther Schmuelling, Carole-Jean Wu, Brian Anderson, Maximilien Breughe, Mark Charlebois, William Chou, Ramesh Chukka, Cody Coleman, Sam Davis, Pan Deng, Greg Diamos, Jared Duke, Dave Fick, J.
+Scott Gardner, Itay Hubara, Sachin Idgunji, Thomas B.
+Jablin, Jeff Jiao, Tom St.
+John, Pankaj Kanwar, David Lee, Jeffery Liao, Anton Lokhmotov, Francisco Massa, Peng Meng, Paulius Micikevicius, Colin Osborne, Gennady Pekhimenko, Arun Tejusve Raghunath Rajan, Dilip Sequeira, Ashish Sirasao, Fei Sun, Hanlin Tang, Michael Thomson, Frank Wei, Ephrem Wu, Lingjie Xu, Koichi Yamada, Bing Yu, George Yuan, Aaron Zhong, Peizhao Zhang, and Yuchen Zhou.
+2019.
+MLPerf Inference Benchmark.
+CoRR abs/1911.02549 (2019).
+arXiv:1911.02549 http://arxiv.org/abs/1911.02549
+- Rein et al.
+(2024) David Rein, Betty Li Hou, Asa Cooper Stickland, Jackson Petty, Richard Yuanzhe Pang, Julien Dirani, Julian Michael, and Samuel R.
+Bowman.
+2024.
+GPQA: A Graduate-Level Google-Proof Q&A Benchmark.
+In First Conference on Language Modeling.
+https://openreview.net/forum?id=Ti67584b98
 - Ribeiro et al.
@@
 2016.
-” Why should i trust you?” Explaining the predictions of any classifier.
+" Why should i trust you?" Explaining the predictions of any classifier.
 In Proceedings of the 22nd ACM SIGKDD international conference on knowledge discovery and data mining.
@@
 High-Resolution Image Synthesis with Latent Diffusion Models.
-In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition.
-10684–10695.
+In IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2022, New Orleans, LA, USA, June 18-24, 2022.
+IEEE, 10674–10685.
+https://doi.org/10.1109/CVPR52688.2022.01042
 - Rudin et al.
@@
 Statistic Surveys 16 (2022), 1–85.
-- Saharia et al.
-(2022) Chitwan Saharia, William Chan, Saurabh Saxena, Lala Li, Jay Whang, Emily Denton, Seyed Kamyar Seyed Ghasemipour, Burcu Karagol Ayan, S.
-Sara Mahdavi, Rapha Gontijo Lopes, Tim Salimans, Jonathan Ho, David J Fleet, and Mohammad Norouzi.
+- Ruehle et al.
+(2021) Victor Ruehle, Robert Sim, Sergey Yekhanin, Nishanth Chandran, Melissa Chase, Daniel Jones, Kim Laine, Boris Köpf, Jaime Teevan, Jim Kleewein, et al.
+2021.
+Privacy preserving machine learning: Maintaining confidentiality and preserving trust.
+https://www.microsoft.com/en-us/research/blog/privacy-preserving-machine-learning-maintaining-confidentiality-and-preserving-trust/
+- Sawadogo and Darmont (2021) Pegdwendé Sawadogo and Jérôme Darmont. 2021. On data lake architectures and metadata management. Journal of Intelligent Information Systems 56 (2021), 97–120.
+- Schick et al.
+(2024) Timo Schick, Jane Dwivedi-Yu, Roberto Dessì, Roberta Raileanu, Maria Lomeli, Eric Hambro, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom.
+2024.
+Toolformer: Language models can teach themselves to use tools.
+Advances in Neural Information Processing Systems 36 (2024).
+- Schramowski et al.
+(2023) Patrick Schramowski, Manuel Brack, Björn Deiseroth, and Kristian Kersting.
+2023.
+Safe latent diffusion: Mitigating inappropriate degeneration in diffusion models.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+22522–22531.
+- Schürholt et al.
+(2022) Konstantin Schürholt, Diyar Taskiran, Boris Knyazev, Xavier Giró-i Nieto, and Damian Borth.
 2022.
-Photorealistic Text-to-Image Diffusion Models with Deep Language Understanding.
-arXiv:2205.11487 [cs.CV]
-- Sawadogo and Darmont (2021) Pegdwendé Sawadogo and Jérôme Darmont. 2021. On data lake architectures and metadata management. Journal of Intelligent Information Systems 56 (2021), 97–120.
+Model zoos: A dataset of diverse populations of neural network models.
+Advances in Neural Information Processing Systems 35 (2022), 38134–38148.
+- Schwarzschild et al.
+(2024) Avi Schwarzschild, Zhili Feng, Pratyush Maini, Zachary C.
+Lipton, and J.
+Zico Kolter.
+2024.
+Rethinking LLM Memorization through the Lens of Adversarial Compression.
+arXiv:cs.LG/2404.15146 https://arxiv.org/abs/2404.15146
 - Security ([n.d.]a) Mithril Security. [n.d.]a. AICert – Open-source tool to trace AI models’ provenance. https://www.mithrilsecurity.io/aicert.
@@
 - Security ([n.d.]c) Mithril Security. [n.d.]c. Welcome to AICert! https://aicert.mithrilsecurity.io/en/latest/.
+- Sharkey et al.
+(2025) Lee Sharkey, Bilal Chughtai, Joshua Batson, Jack Lindsey, Jeff Wu, Lucius Bushnaq, Nicholas Goldowsky-Dill, Stefan Heimersheim, Alejandro Ortega, Joseph Bloom, Stella Biderman, Adria Garriga-Alonso, Arthur Conmy, Neel Nanda, Jessica Rumbelow, Martin Wattenberg, Nandi Schoots, Joseph Miller, Eric J.
+Michaud, Stephen Casper, Max Tegmark, William Saunders, David Bau, Eric Todd, Atticus Geiger, Mor Geva, Jesse Hoogland, Daniel Murfet, and Tom McGrath.
+2025.
+Open Problems in Mechanistic Interpretability.
+arXiv:cs.LG/2501.16496 https://arxiv.org/abs/2501.16496
 - Shen et al.
@@
 Advances in Neural Information Processing Systems 36 (2024).
+- Shi et al.
+(2024) Weijia Shi, Anirudh Ajith, Mengzhou Xia, Yangsibo Huang, Daogao Liu, Terra Blevins, Danqi Chen, and Luke Zettlemoyer.
+2024.
+Detecting Pretraining Data from Large Language Models.
+In The Twelfth International Conference on Learning Representations.
+https://openreview.net/forum?id=zWqr3MQuNs
+- Shokri et al.
+(2017) Reza Shokri, Marco Stronati, Congzheng Song, and Vitaly Shmatikov.
+2017.
+Membership Inference Attacks Against Machine Learning Models.
+In 2017 IEEE Symposium on Security and Privacy (SP).
+3–18.
+https://doi.org/10.1109/SP.2017.41
+- Shraga and Miller (2023) Roee Shraga and Renée J.
+Miller.
+2023.
+Explaining Dataset Changes for Semantic Data Versioning with Explain-Da-V.
+Proc.
+VLDB Endow.
+16, 6 (2023), 1587–1600.
+https://doi.org/10.14778/3583140.3583169
 - Singh et al.
@@
 Nature Communications 14, 1 (2023), 7913.
-- Singh et al. (2024) Chandan Singh, Jeevana Priya Inala, Michel Galley, Rich Caruana, and Jianfeng Gao. 2024. Rethinking Interpretability in the Era of Large Language Models. arXiv:2402.01761 [cs.CL]
+- Singh et al. (2024) Chandan Singh, Jeevana Priya Inala, Michel Galley, Rich Caruana, and Jianfeng Gao. 2024. Rethinking Interpretability in the Era of Large Language Models. arXiv:cs.CL/2402.01761
 - Sinitsin et al.
@@
 https://openreview.net/forum?id=HJedXaEtvS
+- Spadari et al.
+(2024) Vincenzo Spadari, Francesco Cerasuolo, Giampaolo Bovenzi, and Antonio Pescapè.
+2024.
+An MLOps Framework for Explainable Network Intrusion Detection with MLflow.
+In 2024 IEEE Symposium on Computers and Communications (ISCC).
+1–6.
+https://doi.org/10.1109/ISCC61673.2024.10733700
+- Staab et al.
+(2025) Robin Staab, Mark Vero, Mislav Balunovic, and Martin Vechev.
+2025.
+Language Models are Advanced Anonymizers.
+In The Thirteenth International Conference on Learning Representations.
+https://openreview.net/forum?id=82p8VHRsaK
 - Stoyanovich and Howe (2019) Julia Stoyanovich and Bill Howe. 2019. Nutritional Labels for Data and Models. IEEE Data Eng. Bull. 42, 3 (2019), 13–23. http://sites.computer.org/debull/A19sept/p13.pdf
+- Sun et al. (2025) Mingjie Sun, Yida Yin, Zhiqiu Xu, J. Zico Kolter, and Zhuang Liu. 2025. Idiosyncrasies in Large Language Models. arXiv:cs.CL/2502.12150 https://arxiv.org/abs/2502.12150
 - Sundararajan et al. (2017) Mukund Sundararajan, Ankur Taly, and Qiqi Yan. 2017. Axiomatic attribution for deep networks. In International conference on machine learning. PMLR, 3319–3328.
-- Touvron et al.
-(2023) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom.
-2023.
-Llama 2: Open Foundation and Fine-Tuned Chat Models.
-arXiv:2307.09288 [cs.CL]
+- Tabassi (2023) Elham Tabassi. 2023. Artificial Intelligence Risk Management Framework (AI RMF 1.0). https://doi.org/10.6028/NIST.AI.100-1
+- Tabassi (2024) Elham Tabassi.
+2024.
+Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 laying down harmonised rules on artificial intelligence and amending Regulations (EC) No 300/2008, (EU) No 167/2013, (EU) No 168/2013, (EU) 2018/858, (EU) 2018/1139 and (EU) 2019/2144 and Directives 2014/90/EU, (EU) 2016/797 and (EU) 2020/1828 (Artificial Intelligence Act) (Text with EEA relevance) PE/24/2024/REV/1, OJ L, 2024/1689, 12.7.2024.
+https://eur-lex.europa.eu/eli/reg/2024/1689/oj
 - Traub et al.
@@
 - Ustun and Rudin (2016) Berk Ustun and Cynthia Rudin. 2016. Supersparse linear integer models for optimized medical scoring systems. Machine Learning 102 (2016), 349–391.
-- Vartak et al.
-(2018) Manasi Vartak, Joana M.
-F.
-da Trindade, Samuel Madden, and Matei Zaharia.
-2018.
-MISTIQUE: A System to Store and Query Model Intermediates for Model Diagnosis.
-In Proceedings of the 2018 International Conference on Management of Data (Houston, TX, USA) (SIGMOD ’18).
-Association for Computing Machinery, New York, NY, USA, 1285–1300.
-https://doi.org/10.1145/3183713.3196934
-- Vaswani et al.
-(2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N.
-Gomez, Łukasz Kaiser, and Illia Polosukhin.
-2017.
-Attention is all you need.
-In Proceedings of the 31st International Conference on Neural Information Processing Systems (Long Beach, California, USA) (NIPS’17).
-Curran Associates Inc., Red Hook, NY, USA, 6000–6010.
 - Verma et al.
@@
 https://aws.amazon.com/blogs/machine-learning/build-a-cross-account-mlops-workflow-using-the-amazon-sagemaker-model-registry/.
+- Vishwambari and Agrawal (2023) T Vishwambari and Sonali Agrawal.
+2023.
+Integration of Open-Source Machine Learning Operations Tools into a Single Framework.
+In 2023 International Conference on Computing, Communication, and Intelligent Systems (ICCCIS).
+335–340.
+https://doi.org/10.1109/ICCCIS60361.2023.10425558
+- Wang et al.
+(2024a) Keyu Wang, Abdullah Norozi Iranzad, Scott Schaffter, Doina Precup, and Jonathan Lebensold.
+2024a.
+Mitigating Downstream Model Risks via Model Provenance.
+arXiv:cs.LG/2410.02230 https://arxiv.org/abs/2410.02230
 - Wang et al.
@@
 In The Eleventh International Conference on Learning Representations.
+- Wang et al.
+(2023) Sheng-Yu Wang, Alexei A Efros, Jun-Yan Zhu, and Richard Zhang.
+2023.
+Evaluating data attribution for text-to-image models.
+In Proceedings of the IEEE/CVF International Conference on Computer Vision.
+7192–7203.
+- Wang et al.
+(2024b) Wenxiao Wang, Weiming Zhuang, and Lingjuan Lyu.
+2024b.
+Towards Fundamentally Scalable Model Selection: Asymptotically Fast Update and Selection.
+CoRR abs/2406.07536 (2024).
+https://doi.org/10.48550/ARXIV.2406.07536 arXiv:2406.07536
+- Wei et al.
+(2022) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H.
+Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus.
+2022.
+Emergent Abilities of Large Language Models.
+Transactions on Machine Learning Research (2022).
+https://openreview.net/forum?id=yzkSU5zdwD Survey Certification.
+- Winata et al.
+(2024) Genta Indra Winata, Hanyang Zhao, Anirban Das, Wenpin Tang, David D.
+Yao, Shi-Xiong Zhang, and Sambit Sahu.
+2024.
+Preference Tuning with Human Feedback on Language, Speech, and Vision Tasks: A Survey.
+arXiv:cs.CL/2409.11564 https://arxiv.org/abs/2409.11564
+- Xin et al.
+(2024) Rui Xin, Niloofar Mireshghallah, Shuyue Stella Li, Michael Duan, Hyunwoo Kim, Yejin Choi, Yulia Tsvetkov, Sewoong Oh, and Pang Wei Koh.
+2024.
+A False Sense of Privacy: Evaluating Textual Data Sanitization Beyond Surface-level Privacy Leakage.
+In Neurips Safe Generative AI Workshop 2024.
+https://openreview.net/forum?id=3JLtuCozOU
+- Yan et al.
+(2024) Biwei Yan, Kun Li, Minghui Xu, Yueyan Dong, Yue Zhang, Zhaochun Ren, and Xiuzhen Cheng.
+2024.
+On protecting the data privacy of large language models (llms): A survey.
+arXiv preprint arXiv:2403.05156 (2024).
+- Youssef et al. (2024) Paul Youssef, Zhixue Zhao, Jörg Schlötterer, and Christin Seifert. 2024. Detecting Edited Knowledge in Language Models. arXiv:cs.CL/2405.02765 https://arxiv.org/abs/2405.02765
 - Zhao et al.
@@
 Analysis-oriented Metadata for Data Lakes.
-In Proceedings of the 25th International Database Engineering & Applications Symposium (Montreal, QC, Canada) (IDEAS ’21).
+In Proceedings of the 25th International Database Engineering & Applications Symposium (IDEAS ’21).
 Association for Computing Machinery, New York, NY, USA, 194–203.
 https://doi.org/10.1145/3472163.3472273
-- Zheng et al.
-(2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric.
-P Xing, Hao Zhang, Joseph E.
-Gonzalez, and Ion Stoica.
-2023.
-Judging LLM-as-a-judge with MT-Bench and Chatbot Arena.
-arXiv:2306.05685 [cs.CL]
-- Zhu et al. (2020) Chen Zhu, Ankit Singh Rawat, Manzil Zaheer, Srinadh Bhojanapalli, Daliang Li, Felix Yu, and Sanjiv Kumar. 2020. Modifying Memories in Transformer Models. arXiv:2012.00363 [cs.CL]
+- Zhou et al.
+(2024) Andy Zhou, Bo Li, and Haohan Wang.
+2024.
+Robust Prompt Optimization for Defending Language Models Against Jailbreaking Attacks.
+In The Thirty-eighth Annual Conference on Neural Information Processing Systems.
+https://openreview.net/forum?id=jXs6Cvpe7k
+- Zhou et al.
+(2025) Zhanpeng Zhou, Zijun Chen, Yilan Chen, Bo Zhang, and Junchi Yan.
+2025.
+On the emergence of cross-task linearity in pretraining-finetuning paradigm.
+In Proceedings of the 41st International Conference on Machine Learning (ICML’24).
+JMLR.org, Article 2559, 31 pages.
 - Zhu et al. (2019) Erkang Zhu, Dong Deng, Fatemeh Nargesian, and Renée J Miller. 2019. Josie: Overlap set similarity search for finding joinable tables in data lakes. In SIGMOD. 847–864.
@@
 https://doi.org/10.14778/2994509.2994534
+- Zhu et al. (2025) Sally Zhu, Ahmed Ahmed, Rohith Kuditipudi, and Percy Liang. 2025. Independence Tests for Language Models. arXiv:cs.LG/2502.12292 https://arxiv.org/abs/2502.12292
 - Zou et al.
-(2023) Andy Zou, Long Phan, Sarah Chen, James Campbell, Phillip Guo, Richard Ren, Alexander Pan, Xuwang Yin, Mantas Mazeika, Ann-Kathrin Dombrowski, et al.
-2023.
+(2023a) Andy Zou, Long Phan, Sarah Chen, James Campbell, Phillip Guo, Richard Ren, Alexander Pan, Xuwang Yin, Mantas Mazeika, Ann-Kathrin Dombrowski, et al.
+2023a.
 Representation engineering: A top-down approach to ai transparency.
 arXiv preprint arXiv:2310.01405 (2023).
+- Zou et al.
+(2023b) Andy Zou, Zifan Wang, Nicholas Carlini, Milad Nasr, J Zico Kolter, and Matt Fredrikson.
+2023b.
+Universal and transferable adversarial attacks on aligned language models.
+arXiv preprint arXiv:2307.15043 (2023).